### v1

# Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ and Noether-Lefschetz theory of $\mathcal{A}_{g}$

Abstract.
We calculate the genus 1 Gromov-Witten theory of the Hilbert scheme $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ of points in the plane. The fundamental 1-point invariant (with a divisor insertion) is calculated using a correspondence with the families local curve Gromov-Witten theory over the moduli space ¯$\overline{\mathcal{M}}_{1,1}$. The answer exactly matches a parallel calculation related to the Noether-Lefschetz geometry of the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties. As a consequence, we prove that the associated cycle classes satisfy a homomorphism property for the projection operator on $\mathsf{CH}^{*}(\mathcal{A}_{g})$. The fundamental 1-point invariant determines the full genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ modulo a nondegeneracy conjecture about the quantum cohomology. A table of calculations is given.
0. Introduction
0.1. Hilbert schemes
The Hilbert scheme $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ of $n$ points in the plane $\mathbb{C}^{2}$ is a nonsingular, irreducible, quasi-projective variety of dimension $2n$ parameterizing ideals $\mathcal{I}\subset\mathbb{C}[x,y]$ of colength $n$,
$\dim_{\mathbb{C}}{\mathbb{C}[x,y]}/{\mathcal{I}}=n\,.$
An open dense set of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ parameterizes ideals associated to configurations of $n$ distinct unordered points. The geometry of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ has been studied from many points of view for several decades now, see [11, 16, 17, 20, 32]. Our perspective here is related to the interactions of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ with Gromov-Witten theory and the relative Donaldson-Thomas invariants of threefolds as developed in [3, 27, 28, 29, 30, 37, 38, 39, 43, 44, 45].
The algebraic torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ acts diagonally on $\mathbb{C}^{2}$ by scaling coordinates,
$(z_{1},z_{2})\cdot(x,y)=(z_{1}x,z_{2}y)\,.$
Let $t_{1}$ and $t_{2}$ denote the equivariant parameters corresponding to the weights of the $\mathsf{T}$-action on the tangent space $\text{Tan}_{0}(\mathbb{C}^{2})$ at the origin of $\mathbb{C}^{2}$.
There is a canonically induced $\mathsf{T}$-action on $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$. The associated $\mathsf{T}$-equivariant cohomology, $H_{\mathsf{T}}^{*}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),{\mathbb{Q}})$, admits a natural basis (as a $\mathbb{Q}[t_{1},t_{2}]$-module) called the Nakajima basis. The Nakajima basis element $\left|\mu\right\rangle$ corresponding to the partition $\mu$ of $n$ is
$\frac{1}{\Pi_{i}\mu_{i}}[V_{\mu}]$
where $[V_{\mu}]$ is (the cohomological dual of) the class of the subvariety of $\mathsf{Hilb}^{|\mu|}(\mathbb{C}^{2})$ with generic element given by a union of schemes of lengths
$\mu_{1},\ldots,\mu_{\ell(\mu)}$
supported at $\ell(\mu)$ distinct points [^1] The points and parts of $\mu$ are considered here to be unordered. of $\mathbb{C}^{2}$. The element $\left|1^{n}\right\rangle$ corresponds to the unit
$\mathsf{1}\in H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),{\mathbb{Q}})\,.$
See [33] for a foundational treatment.
The Hilbert scheme carries a tautological rank $n$ vector bundle,
(0.1) $\mathcal{O}/\mathcal{I}\rightarrow\mathsf{Hilb}^{n}(\mathbb{C}^{2})\,,$
with fiber $\mathbb{C}[x,y]/\mathcal{I}$ over $[\mathcal{I}]\in\mathsf{Hilb}^{n}(\mathbb{C}^{2})$. The $\mathsf{T}$-action on $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ lifts canonically to the tautological bundle (0.1). Let
$D=c_{1}(\mathcal{O}/\mathcal{I})\in H^{2}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),{\mathbb{Q}})$
be the $\mathsf{T}$-equivariant first Chern class. A straightforward calculation [^2] The $n=0,1$ cases are degenerate: $D=0$ for both. for $n\geq 2$ shows
$D=-\left|2,1^{n-2}\right\rangle\,,$
see [25].
The $\mathsf{T}$-equivariant quantum cohomology of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ has been determined in [37]. The matrix elements of the $\mathsf{T}$-equivariant quantum product count [^3] The count is virtual. rational curves meeting three given subvarieties of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$. The (non-negative) degree of an effective [^4] The $\beta=0$ is considered here effective. curve class
$\beta\in H_{2}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),{\mathbb{Z}})$
is defined by pairing with $D$,
$d=\int_{\beta}D\,.$
Curves of degree $d$ are counted with weight $q^{d}$, where $q$ is the quantum parameter. The ordinary multiplication in $\mathsf{T}$-equivariant cohomology is recovered by setting $q=0$.
Let $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ be the operator of quantum multiplication [^5] Here, the symbol $\star$ denotes the small quantum product. In Section 4.3, the large quantum product will also play a role (and will be denoted by $\star_{t}$). by the divisor $D$,
$\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}:QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))\rightarrow QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))\,,\ \ \ \ \ \mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(\gamma)=D\star\gamma\,.$
The operator $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$ is calculated explicitly in the Nakajima basis for all $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ in [37]. The matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$ lie in the field of rational functions in $q$ (with coefficients [^6] In the context of quantum cohomology, the definitions require localization, so we will always consider $H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ and $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as modules over $\mathbb{Q}(t_{1},t_{2})$. in $\mathbb{Q}(t_{1},t_{2})$) and determine all genus 00 Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ by [37, Section 4.2].
We are interested here in the $\mathsf{T}$-equivariant Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ in genus 1. Let $\mu^{1},\ldots,\mu^{r}$ be partitions of $n$. Define
(0.2) $\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\,=\,\sum_{d=0}^{\infty}\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1,d}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}q^{d}\,\in\mathbb{Q}(t_{1},t_{2})[[q]].$
The series (0.2) is always a rational function in $q$,
$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2})(q).\,$
The first nontrivial computation in genus 1 appeared in [44]:
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{2}(\mathbb{C}^{2})}=-\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{q+1}{q-1}\,.$
Our goal here is to provide calculations of all of the series (0.2) starting with the basic case of $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$.
0.2. Gromov-Witten theory of families of local elliptic curves
Let ¯$\overline{\mathcal{M}}_{1,r}$ be the moduli space of Deligne-Mumford stable curves of genus $1$ with $r$ markings. [^7] We will always assume $r>0$ for stability. Let
$\pi:\mathcal{E}\rightarrow\overline{\mathcal{M}}_{1,r}\,$
be the universal elliptic curve with sections
$\mathsf{p}_{1},\ldots,\mathsf{p}_{r}:\overline{\mathcal{M}}_{1,r}\rightarrow\mathcal{E}$
associated to the markings. Let
$\pi_{\mathbb{C}^{2}}:\mathcal{E}\times\mathbb{C}^{2}\rightarrow\overline{\mathcal{M}}_{1,r}$
be the universal local curve over ¯$\overline{\mathcal{M}}_{1,r}$. The torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ acts on the $\mathbb{C}^{2}$ factor as before.
Let $\mu^{1},\ldots,\mu^{r}\in\mathrm{Part}(n)$, and let ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ be the moduli space of stable [^8] The superscript $\bullet$ indicates possibly disconnected domain curves (but no connected component of the domain is contracted to a point). We follow the conventions of [3]. relative maps to the fibers of $\pi_{\mathbb{C}^{2}}$,
$\epsilon:\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})\rightarrow\overline{\mathcal{M}}_{1,r}\,.$
The fiber of $\epsilon$ over the moduli point
$(E,p_{1},\ldots,p_{r})\in\overline{\mathcal{M}}_{1,r}\,$
is the moduli space of stable maps of genus $g$ to $E\times\mathbb{C}^{2}$ relative to the divisors determined by the nodes and the markings of $E$ with boundary [^9] The boundary conditions are unordered, and the cohomology weights of the boundary conditions are all the identity class. condition $\mu^{i}$ over the divisor $p_{i}\times\mathbb{C}^{2}$. Since the degree $n$ is recorded in the size of the partitions $\mu^{i}$, we omit $n$ from the notation for ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$.
The moduli space ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ has $\pi_{\mathbb{C}^{2}}$-relative virtual dimension
$-nr+\sum_{i=1}^{r}\ell(\mu^{i})\,.$
The Gromov-Witten series of the family $\pi_{\mathbb{C}^{2}}$ is defined by
$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}^{\pi_{\mathbb{C}^{2}},\bullet}=\sum_{b=0}^{\infty}{u^{b}}\int_{\overline{\mathcal{M}}_{1,r}}\epsilon_{*}\left(\Big{[}\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}\right).$
Here, the summation index $b$ is the branch point number, so
$2g[b]-2=b+nr-\sum_{i=1}^{r}\ell(\mu^{i})\,.$
The moduli space of stable maps ¯$\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}$ is empty unless $g[b]$ is an integer. The virtual class [¯$\Big{[}\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}$ is the $\pi_{\mathbb{C}^{2}}$-relative $\mathsf{T}$-equivariant virtual class of the family of relative stable maps to the fibers of $\pi_{\mathbb{C}^{2}}$. We define
$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{g}^{\pi_{\mathbb{C}^{2}},\bullet}$ $=$ $\text{Coeff}_{u^{2g-2}}\Big{[}\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}^{\pi_{\mathbb{C}^{2}},\bullet}\Big{]}$
$=$ ¯$\int_{\overline{\mathcal{M}}_{1,r}}\epsilon_{*}\left(\Big{[}\overline{\mathcal{M}}^{\bullet}_{g}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}\right).$
To emphasize the degree $n$, we will sometimes use the notation
$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{g,n}^{\pi_{\mathbb{C}^{2}},\bullet}=\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{g}^{\pi_{\mathbb{C}^{2}},\bullet}\,.$
The Gromov-Witten series of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ and the Gromov-Witten series of the family $\pi_{\mathbb{C}^{2}}$ are related by the following result of [44].
Theorem A (Pandharipande-Tseng). For all $\mu^{1},\mu^{2},\ldots,\mu^{r}\in\text{\em Part}(n)$, we have
$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=(-i)^{\sum_{i=1}^{r}\ell(\mu^{i})-|\mu^{i}|}\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}^{\pi_{\mathbb{C}^{2}},\bullet}$
after the variable change $-q=e^{iu}$.
The calculation of the Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ in genus 1 is therefore equivalent to the calculation of the Gromov-Witten invariants of the family $\pi_{\mathbb{C}^{2}}$ of local elliptic curves. For example,
(0.3) $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=-(-i)^{-1}\big{\langle}(2,1^{n-2})\big{\rangle}^{\pi_{\mathbb{C}^{2}},\bullet}\,.$
0.3. Moduli of abelian varieties
Let $\mathcal{A}_{g}$ be the moduli space of principally polarized abelian varieties $(X,\theta)$ of dimension $g$. The space $\mathcal{A}_{g}$ is a nonsingular Deligne-Mumford stack of dimension $\binom{g+1}{2}$. Let
$\nu:\mathcal{X}_{g}\rightarrow\mathcal{A}_{g}$
be the universal principally polarized abelian variety. We refer the reader to [2] for the foundations of the study of the moduli of abelian varieties.
A general abelian variety $(X,\theta)$ parameterized by $\mathcal{A}_{g}$ has Picard number $1$. The Noether-Lefschetz locus of $\mathcal{A}_{g}$ parameterizes abelian varieties with Picard number at least 2. The simplest components of the Noether-Lefschetz locus of $\mathcal{A}_{g}$ are related to the geometry of elliptic curves on abelian varieties. For $g\geq 1$ and $n\geq 1$, let
$\mathsf{NL}_{g,n}=\left\{(X,\theta)\in\mathcal{A}_{g}\left|\begin{array}[]{c}X\text{ contains a subgroup }E\subset X\\ \text{ which is an elliptic curve of degree }\theta\cdot[E]=n\end{array}\right.\right\}.$
Let $[{\mathsf{NL}}_{g,n}]\in\mathsf{CH}^{g-1}(\mathcal{A}_{g})$ be the associated cycle class.101010 The $g=1$ case is degenerate: $[\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$ by definition. All Chow groups are taken with $\mathbb{Q}$-coefficients. The following linear combination of components plays a geometrically important role [^11] As usual, $\sigma_{r}(n)=\sum_{d|n}d^{r}$.
$[\widetilde{\mathsf{NL}}_{g,n}]=\sum_{n^{\prime}|n}\sigma_{1}\left(\frac{n}{n^{\prime}}\right)[\mathsf{NL}_{g,n^{\prime}}]\ \in\ \mathsf{CH}^{g-1}(\mathcal{A}_{g})\,,$
see [19].
The Hodge bundle is the rank $g$ vector bundle
$\mathbb{E}_{g}=\nu_{*}(\Omega_{\nu}).$
The Chern classes $\lambda_{i}=c_{i}(\mathbb{E}_{g})\in\mathsf{CH}^{i}(\mathcal{A}_{g})$ generate the tautological ring [48].
$\mathsf{R}^{*}(\mathcal{A}_{g})\subset\mathsf{CH}^{*}(\mathcal{A}_{g})\,.$
A canonical $\mathbb{Q}$-linear projection operator
$\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{*}(\mathcal{A}_{g})$
has been constructed in [4].
The following result of [22] determines the projections of the simplest components of the Noether-Lefschetz loci.
Theorem B (Iribar-López). For $g\geq 2$ and $n\geq 1$, we have
$\mathsf{taut}([\mathsf{NL}_{g,n}])=\frac{n^{2g-1}g}{6|B_{2g}|}\prod_{p|n}(1-p^{2-2g})\lambda_{g-1}$
or, equivalently [^12] Equation (0.4) is correct also for $g=1$ with the convention $[\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$.,
(0.4) $\mathsf{taut}\left(\frac{(-1)^{g}}{24}\lambda_{g-1}+\sum_{n=1}^{\infty}[\widetilde{\mathsf{NL}}_{g,n}]\,Q^{n}\right)=\frac{(-1)^{g}}{24}E_{2g}(Q)\lambda_{g-1},$
where $E_{2g}(Q)$ is the Eisenstein modular function of weight $2g$ in the variable $Q=e^{2\pi i\tau}$.
A basic open question regarding the structure of the projection operator is whether $\mathsf{taut}$ is a homomorphism of $\mathbb{Q}$-algebras:
$\mathsf{taut}(\gamma)\cdot\mathsf{taut}(\widehat{\gamma})\stackrel{{\scriptstyle?}}{{=}}\mathsf{taut}(\gamma\cdot\widehat{\gamma})\ \in\ \mathsf{R}^{*}(\mathcal{A}_{g})$
for all $\gamma,\widehat{\gamma}\in\mathsf{CH}^{*}(\mathcal{A}_{g})$.
0.4. A triple equivalence
0.4.1. Hilbert schemes of points
Let $\mathsf{Tr}_{n}$ be the normalized trace of the operator $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ of quantum multiplication by the divisor $D$ on $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$,
$\mathsf{Tr}_{n}=\frac{1}{t_{1}+t_{2}}\mathsf{trace}(\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})\,.$
Using the formula of [37] for $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$, we obtain
(0.5) $\mathsf{Tr}_{n}=\sum_{\mu\in\mathsf{Part}(n)}\sum_{i}\left(\frac{\mu_{i}^{2}}{2}\frac{(-q)^{\mu_{i}}+1}{(-q)^{\mu_{i}}-1}-\frac{\mu_{i}}{2}\frac{(-q)+1}{(-q)-1}\right)\in\mathbb{Q}(q),$
where the first summation is over the set $\mathsf{Part}(n)$ of partitions of $n$ and the second summation is over all parts $\mu_{i}$ of the partition $\mu\in\mathsf{Part}(n)$.
The first of three equivalent statements is the following calculation of the basic genus 1 Gromov-Witten invariant of the Hilbert scheme of points.
Theorem 1.
For $n\geq 1$, we have
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=-\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\left(\mathsf{Tr}_{n}+\sum_{k=2}^{n-1}{\sigma_{-1}(n-k)}\mathsf{Tr}_{k}\right)\,.$
The evaluations for $n=1,2$ agree with the known results:
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{1}(\mathbb{C}^{2})}=-\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\mathsf{Tr}_{1}=0\,,$
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{2}(\mathbb{C}^{2})}=-\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\mathsf{Tr}_{2}=-\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{q+1}{q-1}\,.$
For $n=3,4,5$, we obtain
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{3}(\mathbb{C}^{2})}$ $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot(\mathsf{Tr}_{3}+\mathsf{Tr}_{2})$
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\frac{5q^{3}-3q^{2}-3q+5}{(q-1)(q^{2}-q+1)}\right)$
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(-5-7q-q^{2}+2q^{3}-q^{4}-7q^{5}-10q^{6}-7q^{7}+...\right),$
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}$ $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\mathsf{Tr}_{4}+\mathsf{Tr}_{3}+\frac{3}{2}\mathsf{Tr}_{2}\right)$
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(\frac{35q^{5}-28q^{4}+23q^{3}+23q^{2}-28q+35}{2(q-1)(q^{2}+1)(q^{2}-q+1)}\right)}}$
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(-\frac{35}{2}-21q-q^{2}-3q^{3}-17q^{4}-21q^{5}-19q^{6}-21q^{7}+...\right)\,,$
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}$ $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\mathsf{Tr}_{5}+\mathsf{Tr}_{4}+\frac{3}{2}\mathsf{Tr}_{3}+\frac{4}{3}\mathsf{Tr}_{2}\right)$
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(-\frac{272q^{9}-539q^{8}+760q^{7}-629q^{6}+302q^{5}+302q^{4}-629q^{3}+760q^{2}-539q+272}{6(q-1)(q^{2}+1)(q^{2}-q+1)(q^{4}-q^{3}+q^{2}-q+1)}\right)}}$
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\frac{136}{3}+\frac{277}{6}q-\frac{41}{6}q^{2}+\frac{17}{3}q^{3}+\frac{151}{6}q^{4}+\frac{127}{6}q^{5}+\frac{101}{3}q^{6}+\frac{277}{6}q^{7}+...\right).$
0.4.2. Families of local elliptic curves
It is natural to look for an analogue of Theorem 1 for the Gromov-Witten theory of families of local elliptic curves using the correspondence of Theorem A. The best statement is expressed in terms of the descendent Gromov-Witten theory of the family
$\pi:\mathcal{E}\rightarrow\overline{\mathcal{M}}_{1,1}\,$
for stable maps with connected domains.
Let ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ be the moduli space of stable [^13] The superscript $\circ$ indicates connected domain curves. relative maps to the fibers of $\pi$,
$\epsilon:\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)\rightarrow\overline{\mathcal{M}}_{1,1}\,.$
The fiber of $\epsilon$ over the moduli point
$(E,p_{1})\in\overline{\mathcal{M}}_{1,1}\,$
is the moduli space of stable maps of 1-pointed connected genus $g$ curves to $E$ of degree $n$ relative to the divisors determined by the nodes of $C$. The moduli space ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ has $\pi$-relative virtual dimension $12g-$ 1 and total virtual dimension $2g$.
For $g\geq 2$, let
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\big{\rangle}_{g,n}^{\pi,\circ}=\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir_{\pi}}}\tau_{1}(\mathsf{p_{1}})\lambda_{g}\lambda_{g-2}\,\in\,\mathbb{Q}\,,$
where $\mathsf{p}_{1}$ is viewed as a divisor class on $\mathcal{E}$ and the Hodge classes $\lambda_{g}\lambda_{g-2}$ are pulled-back from the moduli of the domain curves. [^14] The $\lambda_{g}$ class plays an important role in the theory Hodge integrals, see [10, 13, 31].
Theorem 2.
For $g\geq 2$, we have
$\sum_{n=0}^{\infty}\big{\langle}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\big{\rangle}_{g,n}^{\pi,\circ}Q^{n}=\frac{(-1)^{g}}{24}\frac{|B_{2g}|}{4g}\frac{|B_{2g-2}|}{(2g-2)!}E_{2g}(Q)\,.$
The $g=1$ case is degenerate. The corresponding evaluation is
(0.6) $\sum_{n=0}^{\infty}\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}Q^{n}=\frac{(-1)}{24}\frac{|B_{2}|}{4}\frac{|B_{0}|}{0!}E_{2}(Q)=-\frac{1}{576}E_{2}(Q)\,.$
Theorem A together with a families relative/descendent correspondence will be used to show that Theorem 1 and 2 are equivalent.
0.4.3. The homomorphism question
The third equivalence involves the homomorphism question for the projection operator
$\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{*}(\mathcal{A}_{g})\,.$
Let $\mathcal{M}_{g}^{\mathsf{ct}}$ be the $33g-$ 3 dimensional moduli space of genus $g$ curves of compact type, and let
$\mathsf{Tor}:\mathcal{M}_{g}^{\mathsf{ct}}\rightarrow\mathcal{A}_{g}$
be the Torelli map. Since $\mathsf{Tor}$ is proper, we can push forward the fundamental class [^15] The $g=1$ case is degenerate: $\mathsf{Tor}_{*}[\mathcal{M}_{1}^{\mathsf{ct}}]=[\mathcal{A}_{1}]$ by definition.:
$\mathsf{Tor}_{*}[\mathcal{M}_{g}^{\mathsf{ct}}]\in\mathsf{CH}^{{\binom{g+1}{2}}-3g+3}(\mathcal{A}_{g})\,.$
The intersection theory of the Torelli cycle has been studied in [5, 7, 22].
Theorem 3.
For $g\geq 1$ and $n\geq 1$, we have
$\mathsf{taut}(\mathsf{Tor}_{*}[\mathcal{M}_{g}^{\mathsf{ct}}])\cdot\mathsf{taut}([\mathsf{NL}_{g,n}])\,=\,\mathsf{taut}(\mathsf{Tor}_{*}[\mathcal{M}_{g}^{\mathsf{ct}}]\cdot[\mathsf{NL}_{g,n}])\ \in\ \mathsf{R}^{*}(\mathcal{A}_{g})\,.$
In other words, the classes $\mathsf{Tor}_{*}[\mathcal{M}_{g}^{\mathsf{ct}}],[\mathsf{NL}_{g,n}]\in\mathsf{CH}^{*}(\mathcal{A}_{g})$ satisfy the multiplicative property with respect to $\mathsf{taut}$. Theorem B together with a study of the geometry of the Torelli map was used to show that Theorems 2 and 3 are equivalent in [22].
0.4.4. Proof strategy
A central result of the paper is the proof of Theorem 2. As a consequence of the equivalences, Theorems 1 and 3 will then also be proven.
While Gromov-Witten theory is well-developed for the study of a fixed target variety, the main difficulty in proving Theorem 2 is that there are very few techniques which are useful for the calculations of Gromov-Witten invariants in families. Our proof of Theorem 2 uses a mix of Hodge integrals, formulas for the Gromov-Witten theory of a fixed elliptic target, and new constraints on family invariants. The methods yields calculations of more general families Hodge integrals in Section 2.6.
0.5. Reduction of multi-point series to 1-point series
By [37, Section 4.2], the quantum powers of $D$ generate the quantum cohomology $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$. More precisely, the set
(0.7) $\{1,D,D^{*2},D^{*3},\ldots,D^{*(|\mathsf{Part}(n)|-1)}\}$
is a basis of $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as a $\mathbb{Q}(t_{1},t_{2})(q)$-vector space. Our first reduction result is that the genus 1 Gromov-Witten theory (0.2) of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ can be effectively reduced to 1-point series in the basis (0.7).
Theorem 4.
To every genus $1$ series $\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\big{\rangle}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$, there are canonically associated functions
[…]
for which the following equation holds:
$\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\sum_{k=0}^{|\mathsf{Part}(n)|-1}\,\sum_{m=0}^{\ell-1}\,C_{k,m}\cdot\left(q\frac{d}{dq}\right)^{m}\big{\langle}D^{*k}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\,.$
The functions (0.8) are constructed as rational functions of the matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$. The proof of Theorem 4 uses the WDVV relations in genus 0 and Getzler’s relation in genus 1.
0.6. Reduction to $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$
The reduction of the genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to the basic series $\langle D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ is more subtle. Since the quantum cohomology of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is semisimple, the Givental-Teleman reconstruction result [15, 47] reduces the full higher genus Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to the genus 0 theory together with the calculation of the $\mathsf{R}$-matrix. The $\mathsf{R}$-matrix was proven to be determined [^16] Since the Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is equivariant, the associated $\mathsf{CohFT}$ is not conformal. by constant map invariants in [44]. The resulting control of the full Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ was used in [44] to prove both the crepant resolution and the GW/DT correspondences associated to the geometry, but was not sufficient to obtain closed form evaluations of any higher genus series (because of the difficulty in the determination and the lack of closed forms for the $\mathsf{R}$-matrix).
In genus 1, the Givental-Teleman reconstruction formula reduces to an earlier equation due to Givental [14] on the associated Frobenius manifold:
(0.9) $d\mathcal{F}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\frac{1}{2}\sum_{i=1}^{|\mathsf{Part}(n)|}\mathsf{R}_{ii}\,du_{i}+\frac{1}{48}\sum_{i=1}^{|\mathsf{Part}(n)|}d\log\Delta_{i}\,.$
Here, $\mathcal{F}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is the $g=1$ Gromov-Witten potential of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$, the $u_{i}$ are the canonical coordinates, and
$\Delta_{i}=\left\langle\frac{\partial}{\partial u_{i}},\frac{\partial}{\partial u_{i}}\right\rangle^{-1}\,.$
From Theorem 1, we know explicitly the functions
(0.10) $\big{\langle}\underbrace{D,\ldots,D}_{\ell}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\left(q\frac{d}{dq}\right)^{\ell-1}\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$
for $\ell=1,\ldots,|\mathsf{Part}(n)|$. We construct a $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|$ matrix which expresses the functions (0.10) in terms of the functions $\mathsf{R}_{ii}\,\big{|}_{\mathsf{t}=0}$, where $\mathsf{t}$ is the coordinate on the Frobenius manifold. The non-degeneracy of the Wronskian
$\mathsf{W}=\left(\begin{array}[]{cccc}\nabla_{D}u_{1}\,\big{|}_{\mathsf{t}=0}&\nabla_{D}u_{2}\,\big{|}_{\mathsf{t}=0}&\ldots&\nabla_{D}u_{|\mathsf{Part}(n)|}\,|_{\mathsf{t}=0}\\[8.0pt] q\frac{d}{dq}\nabla_{D}u_{1}\,\big{|}_{\mathsf{t}=0}&q\frac{d}{dq}\nabla_{D}u_{2}\,\big{|}_{\mathsf{t}=0}&\ldots&q\frac{d}{dq}\nabla_{D}u_{|\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}\\[8.0pt] \vdots&\vdots&\vdots&\vdots\\[8.0pt] (q\frac{d}{dq})^{|\mathsf{Part}(n)|-1}\nabla_{D}u_{1}\,\big{|}_{\mathsf{t}=0}&(q\frac{d}{dq})^{|\mathsf{Part}(n)|-1}\nabla_{D}u_{2}\,\big{|}_{\mathsf{t}=0}&\ldots&(q\frac{d}{dq})^{|\mathsf{Part}(n)|-1}\nabla_{D}u_{|\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}\end{array}\right)$
implies the invertibility of the system.
Theorem 5.
If $\det(\mathsf{W})\neq 0$, the diagonal entries $\mathsf{R}_{ii}\,\big{|}_{\mathsf{t}=0}$ are determined by $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$.
By Givental’s equation (0.9), the entire genus $1$ Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is then determined by $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ together with genus 0 data. We have checked the non-degeneracy of the Wronskian for $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ for $n\leq 7$ and conjecture the nondegeneracy for all $n$.
Theorem 6.
If $\det(\mathsf{W})\neq 0$, the full genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ can be effectively reconstructed from $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ and $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$.
0.7. Acknowledgements
Conversations with P. Bousseau, S. Canning, L. Drakengren, J. Feusi, F. Greer, C. Lian, S. Molcho, G. Oberdieck, A. Okounkov, D. Oprea, A. Pixton, D. Ranganathan, and J. Schmitt have played an important role in our work. We thank J. Hu and Z. Qin for discussions about their paper [21].
A.I.L. was supported by SNF-200020-219369. R.P. was supported by SNF-200020-219369 and SwissMAP. The first presentation of the results proven here was given by R.P. at the Belgian-Dutch Algebraic Geometry Seminar in Leiden in October 2023 with very helpful discussions with C. Faber, D. Holmes, and B. Moonen afterwards. H.-H. T. was supported in part by a Simons foundation collaboration grant. The project was advanced during several visits of H.-H.T. to the Forschungsinstitut für Mathematik at ETH Zürich.
1. Degree 0 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$
We expand here explicitly the formula of Theorem 1 for the degree 00 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to show the connections to Hilbert scheme calculations of Carlsson and Okounkov [6].
Consider first the generating series over the Hilbert schemes of points:
$\sum_{n=0}^{\infty}\langle D\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\,Q^{n}=\langle D\rangle_{1}^{\mathsf{Hilb}^{2}(\mathbb{C}^{2})}\,Q^{2}+\langle D\rangle_{1}^{\mathsf{Hilb}^{3}(\mathbb{C}^{2})}\,Q^{3}+\langle D\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}\,Q^{4}+...\,,$
where the $n=0,1$ terms vanish since $D=0$ for $n=0,1$. Theorem 1 can be written as
(1.1) $\sum_{n=0}^{\infty}\langle D\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\,Q^{n}=-\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(1+\sum_{n=1}^{\infty}\sigma_{-1}(n)\,Q^{n}\right)\cdot\left(\sum_{n\geq 2}^{\infty}\mathsf{Tr}_{n}\,Q^{n}\right)\,.$
The only $q$-dependence on the right side of (1.1) is via $\mathsf{Tr}_{n}$. After restricting (0.5) to $q=0$, we obtain
(1.2) $\sum_{n\geq 2}^{\infty}\mathsf{Tr}_{n}|_{q=0}\,Q^{n}=-\sum_{n=1}^{\infty}Q^{n}\sum_{\mu\in\mathsf{Part}(n)}\sum_{i}\left(\frac{\mu_{i}^{2}-\mu_{i}}{2}\right).$
Let $\mathcal{P}(Q)=\prod_{k\geq 1}\frac{1}{1-Q^{k}}$, $\mathcal{E}_{2}(Q)=\sum_{k\geq 1}\frac{kQ^{k}}{1-Q^{k}}$, and $\mathcal{E}_{3}(Q)=\sum_{k\geq 1}\frac{k^{2}Q^{k}}{1-Q^{k}}$. The following identities follow easily from the generating function of $\mathrm{T}(n,a)$, the number of times the part $a$ occurs in all partitions of $n$, see [40]:
$\sum_{k=1}^{\infty}Q^{k}\Bigg{(}\sum_{\mu\in\mathsf{Part}(k)}\sum_{i}\mu_{i}\Bigg{)}$ $=$ $\mathcal{P}(Q)\cdot\mathcal{E}_{2}(Q)\,,$
$\sum_{k=1}^{\infty}Q^{k}\Bigg{(}\sum_{\mu\in\mathsf{Part}(k)}\sum_{i}\mu_{i}^{2}\Bigg{)}$ $=$ $\mathcal{P}(Q)\cdot\mathcal{E}_{3}(Q)\,.$
After combining with (1.2), we obtain
$\sum_{n=0}^{\infty}\mathsf{Tr}_{n}|_{q=0}\,Q^{n}=\frac{1}{2}\mathcal{P}(Q)(\mathcal{E}_{2}(Q)-\mathcal{E}_{3}(Q))\,.$
The geometric formula for the $q=0$ term of $\langle D\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is obtained via identification of the virtual fundamental class:
[¯$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(D)$ $=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
$=$ $-\frac{1}{24}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})\,.$
Therefore, the $q=0$ part of Theorem 1 is equivalent to the following identity of generating functions:
(1.3) $\sum_{n=0}^{\infty}Q^{n}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})=\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}(1+\log\mathcal{P}(Q))\cdot\frac{1}{2}\mathcal{P}(Q)(\mathcal{E}_{2}(Q)-\mathcal{E}_{3}(Q))\,.$
Consider now the series $\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}$ of [6, Corollary 3] where we take the line bundle $\mathcal{L}$ to be $\mathcal{O}$ with equivariant weight $m$ as in [6, Section 2.1.1]:
$\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}$ $=$ $\sum_{n=0}^{\infty}Q^{n}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})},m)$
$=$ $\sum_{n=0}^{\infty}Q^{n}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot\left(c_{\mathsf{top}}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})+m\cdot c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})+m^{2}\cdot c_{\mathsf{top}-2}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})\ldots\right)\,.$
The left side of (1.3) is the coefficient of $m^{1}$ of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$. By [6, Corollary 3] together with evaluations from [6, Section 2.2.2], we have
$\frac{\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}}{\big{\langle}1\big{\rangle}}=\frac{1}{2}(\mathcal{E}_{2}(Q)-\mathcal{E}_{3}(Q))\cdot\frac{(t_{1}+t_{2})(t_{1}+m)(t_{2}+m)}{t_{1}t_{2}}\,,$
where the series $\big{\langle}1\big{\rangle}$ is
$\big{\langle}1\big{\rangle}$ $=$ $\sum_{n=0}^{\infty}Q^{n}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})},m)$
$=$ $\prod_{n\geq 1}(1-Q^{n})^{(\frac{m(-t_{1}-t_{2}-m)}{t_{1}t_{2}}-1)}\,$
by [6, Corollary 1]. The matching (1.3) then follows from a simple algebraic expansion of the $m^{1}$ coefficient of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$.
The degree 0 term of Gromov-Witten series $\big{\langle}1\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ can be studied similarly. The geometric formula for the $q=0$ term of $\langle 1\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is
[¯$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(1)$ $=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
$=$ $-\frac{1}{24}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})\,,$
which equals the $m^{1}$ coefficient of $-\frac{1}{24}\big{\langle}1\big{\rangle}$. A calculation then yields
$\text{Coeff}_{m^{1}}\Big{[}\langle 1\rangle\Big{]}=\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\mathcal{P}(Q)\log\mathcal{P}(Q)\,.$
We obtain the evaluation
$\sum_{n=0}^{\infty}\langle 1\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}|_{q=0}\,Q^{n}=-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\mathcal{P}(Q)\log\mathcal{P}(Q)\,.$
The Gromov-Witten series $\langle 1\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is much simpler than $\langle D\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$. Because of the axiom of the fundamental class, all positive degree terms of $\langle 1\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ vanish. Therefore,
$\langle 1\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\text{Coeff}_{Q^{n}}\Big{[}-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\mathcal{P}(Q)\log\mathcal{P}(Q)\Big{]}\,.$
2. Families Hodge integrals
2.1. Overview
Our first result is the calculation of the families Hodge integral of Theorem 2 over the moduli spaces of stable maps with connected domains,
$\epsilon:\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)\rightarrow\overline{\mathcal{M}}_{1,1}\,,\ \ \ \pi:\mathcal{E}\rightarrow\overline{\mathcal{M}}_{1,1}\,.$
After evaluation of the special $n=1$ and $g=1$ cases by hand, the main motivation is to use the vanishing of the virtual class of the moduli space of stable maps to a $K3$ surface to prove Theorem 2.
2.2. The $n=0$ case
The $n=0$ invariant concerns the moduli space of degree 0 stable maps ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)$. After imposing the evaluation condition on the marking, the moduli space is
¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)\supset\text{ev}_{1}^{-1}(\mathsf{p}_{1})=\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}\,$
with obstruction bundle $\mathbb{E}_{g}^{\vee}\otimes\mathsf{T}$, where $\mathsf{T}$ is the tangent line on ¯$\overline{\mathcal{M}}_{1,1}$ associated to the marking.
$\bullet$ For $g\geq 2$, we evaluate the $n=0$ terms by:
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\big{\rangle}_{g,0}^{\pi,\circ}$ $=$ [¯$\int_{[\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}\lambda_{g}\lambda_{g-2}$
$=$ ¯$\int_{\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\lambda_{g}\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{T})$
$=$ $(-1)^{g}\int_{\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\lambda_{g}\lambda_{g-2}(\lambda_{g}+\lambda_{g-1}\boxtimes\psi)$
$=$ $\frac{(-1)^{g}(2g-2)}{24}\frac{|B_{2g}||B_{2g-2}|}{4g(2g-2)(2g-2)!}$
$=$ $\frac{(-1)^{g}}{24}\frac{|B_{2g}|}{4g}\frac{|B_{2g-2}|}{(2g-2)!}\,.$
Here, $\psi_{1}$ denotes the cotangent line at the marking of ¯$\overline{\mathcal{M}}_{g,1}$. In the third equality, we have denoted the dual of $c_{1}(\mathsf{T})$ by $\psi$, the cotangent line at the marking of ¯$\overline{\mathcal{M}}_{1,1}$. In the fourth equality, we have used the dilaton equation and the Hodge integral evaluated in [9, Theorem 4]. The final evaluation agrees with Theorem 2 since the Eisenstein series start with 1,
(2.1) $E_{2g}(Q)=1-\frac{4g}{B_{2g}}\sum_{n=1}^{\infty}\sigma_{2g-1}(n)Q^{n}\,.$
$\bullet$ For $g=1$, we evaluate the $n=0$ term by:
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,0}^{\pi,\circ}$ $=$ [¯$\int_{[\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}$
$=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\cdot\mathsf{e}(\mathbb{E}_{1}^{\vee}\otimes\mathsf{T})$
$=$ $(-1)\int_{\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}(\lambda_{1}+\lambda_{0}\boxtimes\psi)$
$=$ $-\frac{1}{576}\,,$
which agrees with the constant term on the right side of (0.6).
2.3. The $g=1$ case
For $g=1$ and $n\geq 1$, we can evaluate the integral $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ by hand. There are no branch points, so the cotangent line on the domain is pulled-back from the cotangent line of ¯$\overline{\mathcal{M}}_{1,1}$. Using the latter cotangent line class, we can express $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ as $\frac{1}{24}$ times a Gromov-Witten invariant of maps to a fixed elliptic target $(E,p_{1})$:
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}=\frac{1}{24}\big{\langle}\tau_{0}({p}_{1})\big{\rangle}_{1,n}^{E,\circ}\,.$
Using the well-known evaluation
$\sum_{n=0}^{\infty}\big{\langle}\tau_{0}({p}_{1})\big{\rangle}_{1,n}^{E,\circ}=-\frac{1}{24}E_{2}(Q)\,,$
we deduce (0.6) for all [^17] The $n=0$ term $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,0}^{\pi,\circ}$ also matches $\frac{1}{24}\big{\langle}\tau_{0}({p}_{1})\big{\rangle}_{1,0}^{E,\circ}$. $n\geq 1$.
2.4. Proof of Theorem 2 for $g\geq 2$ and $n>0$
Let $S$ be an elliptically fibered $K3$ surface $S$ with a section181818 The section is denoted by $\mathsf{p}_{1}$ following the conventions related to families of points curves. The class of $\mathsf{p_{1}}$ is a divisor class in the total space $S$ of the family. $\mathsf{p}_{1}$,
${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{S}}$${\mathbb{P}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\,,}$$\scriptstyle{\mathsf{p}_{1}}$
and $24$ nodal fibers $R_{1},\ldots,R_{24}\subset S$. The fibers of $\pi_{S}$ are $1$-pointed genus $1$ stable curves. The map
$\mathbb{P}^{1}\longrightarrow\overline{\mathcal{M}}_{1,1}$
induced by $\pi_{S}$ is of degree $48$. Therefore,
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{48}\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}.$
The moduli space of stable maps to the fibers of $\pi_{S}:S\rightarrow\mathbb{P}^{1}$ lies over $\mathbb{P}^{1}$,
$\epsilon_{S}:\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\rightarrow\mathbb{P}^{1}\,.$
Proposition 7.
The following vanishing holds for $g\geq 2$ and $n>0$:
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)=0\,.$
The motivation for Proposition 7 comes from the vanishing of Gromov-Witten invariants for $K3$ surfaces in non-zero curve classes. The Euler class $\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)$ in the integrand relates the families virtual class for $\pi_{S}$ to the virtual class for maps to $S$. A proof using deformation to the normal cones of the nodal fibers $R_{i}\subset S$ requires a subtle study of the logarithmic degeneration formula. We will prove the vanishing of Proposition 7 via another path in Section 2.5.
Proposition 7 allow us to exchange the families Hodge integral
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}$
for a fixed target Gromov-Witten invariant. We first expand the Euler class as
(2.2) $\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)=(-1)^{g}\lambda_{g}+(-1)^{g-1}\lambda_{g-1}\cdot\epsilon_{S}^{*}(2[\mathsf{pt}])\,,$
where $[\mathsf{pt}]\in\mathsf{CH}_{0}(\mathbb{P}^{1})$ is the point class. By Proposition 7, we obtain
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\Big{(}(-1)^{g}\lambda_{g}+(-1)^{g-1}\lambda_{g-1}\cdot 2\epsilon_{S}^{*}([\mathsf{pt}]))\Big{)}=0\,.$
Hence,
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g}\lambda_{g-2}$ $=$ $2\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-1}\lambda_{g-2}\cdot[\mathsf{pt}]$
$=$ $2\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(E,n)]^{vir}}\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}\,,$
where the last integral is the Gromov-Witten invariant with a fixed elliptic curve target $(E,p_{1})$.
The evaluation of the required integral for $(E,p_{1})$ follows from the methods of [34],
(2.3) [¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(E,n)]^{vir}}\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}=\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,,$
as we will explain in Section 2.6. The integral can also be obtained using the study of the Gromov-Witten theory of target curves [35] and was first calculated by Pixton in [46].
Theorem 2 then follows from
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{24}\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,$
and the definition of the Eisenstein series (2.1). ∎
2.5. Proof of Proposition 7
The moduli space of stable maps to the fibers of $\pi_{S}:S\rightarrow\mathbb{P}^{1}$ lies over $\mathbb{P}^{1}$,
$\epsilon_{S}:\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\rightarrow\mathbb{P}^{1}\,.$
The universal curve $\mu:\mathcal{C}\rightarrow\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ carries a universal evaluation map
$\mathcal{C}$$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$$f$$\mu$$\nu$
to the universally expanded target
$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\times_{\mathbb{P}^{1}}S$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$h$$\nu$$\pi_{M}$
The target $\mathcal{T}$ is a family of elliptic curves over ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ with possible expansion over the 24 nodal fibers $R_{i}$. The only permitted expansions over $R_{i}$ are simple circuits of rational curves.
The relative dualizing sheaf $\omega_{\pi_{S}}$ of the of elliptic fibration $\pi_{S}:S\rightarrow\mathbb{P}^{1}$ is pulled-back from the base,
$\omega_{\pi_{S}}\cong\pi_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\,.$
Since $g$ is an isomorphism (except for collapsing chains of unstable rational curves to nodes),
$\omega_{\nu}\,\cong\,h^{*}\omega_{\pi_{M}}\,\cong\,\nu^{*}\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\,.$
We therefore obtain $\nu_{*}\omega_{\nu}\cong\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})$.
Consider next the stabilization map
$\mathcal{C}$$\mathcal{C}_{\mathsf{st}}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$\mathsf{st}$$\mu$$\mu_{\mathsf{st}}$
By the geometry of relative stable maps, the contraction $\mathsf{st}$ is an isomorphism (except for collapsing chains of unstable rational curves to nodes). Therefore,
$\omega_{\mu}\cong\mathsf{st}^{*}\omega_{\mu_{\mathsf{st}}},$
and we obtain $\mu_{*}\omega_{\mu}\cong\mathbb{E}_{g}$.
Since the moduli space ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ parametrizes relative maps to the fibers of $\pi_{S}$, we have a pull-back map
(2.4) $\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\cong\nu_{*}\omega_{\nu}\ \stackrel{{\scriptstyle f^{*}}}{{\longrightarrow}}\ \mu_{*}\omega_{\mu}\cong\mathbb{E}_{g}$
over ¯$\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. Since $n>0$, the map (2.4) is injective, so we obtain an exact sequence
(2.5) $0\rightarrow\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\rightarrow\mathbb{E}_{g}\rightarrow\mathbb{F}\rightarrow 0\,,$
where $\mathbb{F}$ is a rank $g-1$ vector bundle on ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$. We therefore have a factorization
[…]
on ¯$\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. For the second equality, the restriction of (2.5) to a fiber of $\epsilon_{S}$ is used to obtain
$c_{g-1}(\mathbb{F})|_{\epsilon_{S}^{-1}(\mathsf{pt})}=\lambda_{g-1}\,.$
The factorization (2.6) of $\lambda_{g}$ implies the vanishing
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{Tan}_{\mathbb{P}^{1}})=0$
by (2.2). ∎
2.6. Hodge integrals and Gromov-Witten theory for a fixed elliptic target
The proof of Proposition 7 yields a stronger statement for $n>0$:
$\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{Tan}_{\mathbb{P}^{1}})\cap[\overline{\mathcal{M}}^{\circ}_{g,r}(\pi_{S},n)]^{vir}=0$
in $\operatorname{\mathsf{CH}}^{*}(\overline{\mathcal{M}}_{g,r})$. If $\mathsf{F}(\lambda)$ is a homogeneous [^19] The class $\lambda_{i}$ has degree $i$. polynomial in Hodge classes $\lambda_{i}$ satisfying
$\sum_{i=1}^{r}k_{i}+\text{deg}(\mathsf{F})=g-1\,.$
we obtain (as in Section 2.4):
(2.7) $\left\langle\prod_{i=1}^{m}\tau_{k_{i}}(\mathsf{p_{1}})\prod_{j=m+1}^{r}\tau_{k_{j}+1}(1)\cdot\lambda_{g}\mathsf{F}(\lambda)\right\rangle_{g,n}^{\pi,\circ}=\frac{1}{24}\left\langle\prod_{i=1}^{m}\tau_{k_{i}}(p_{1})\cdot\prod_{j=m+1}^{r}\tau_{k_{j}+1}(1)\cdot\lambda_{g-1}\mathsf{F}(\lambda)\right\rangle_{g,n}^{E,\circ}\,,$
where the $n=0$ case follows from an application of Mumford’s formula.
The Gromov-Witten Hodge integral of a fixed elliptic curve target $(E,p_{1})$ on the right side can be effectively computed. By [34], we have an equality of cycles
$\sum_{n=0}^{\infty}Q^{n}[\overline{\mathcal{M}}^{\circ}_{g,m}(E,n)]^{vir}\cdot\prod_{i=1}^{m}\tau_{0}(\mathsf{p}_{1})\cdot\lambda_{g-1}=\frac{(-1)^{g}(2g-1)!}{(2g-2+m)!}\lambda_{g}\lambda_{g-1}\sum_{i=1}^{m}\prod_{j\neq i}\psi_{j}\left(Q\frac{d}{dQ}\right)^{m-1}E_{2g}(Q)$
in $H^{*}(\overline{\mathcal{M}}_{g,m})$. Therefore, the right hand side of (2.7) is equal to the coefficient of $Q^{n}$ in
$\frac{(-1)^{g}(2g-1)!}{24(2g+m-2)!}\left(Q\frac{d}{dQ}\right)^{m-1}E_{2g}(Q)\sum_{i=1}^{m}\int_{\overline{\mathcal{M}}_{g,r}}\prod_{j=1}^{r}\psi_{j}^{k_{j}+1-\delta_{ij}}\cdot\mathsf{F}(\lambda)\lambda_{g}\lambda_{g-1}\,.$
The integrals over ¯$\overline{\mathcal{M}}_{g,r}$ can be evaluated effectively via Hodge integral techniques.
$\bullet$ For $\mathsf{F}=1$, we have the exact evaluation
¯$\int_{\overline{\mathcal{M}}_{g,r}}\prod_{j=1}^{r}\psi_{j}^{k_{j}+1-\delta_{ij}}\cdot\lambda_{g}\lambda_{g-1}=\frac{|B_{2g}|}{2^{2g-1}(2g)!}\frac{(2g+r-3)!(2k_{i}+1)}{(2k_{1}+1)!!\ldots(2k_{r}+1)!!}\,,$
given by the Virasoro constraints of $\mathbb{P}^{2}$ [13]. We arrive at the following:
$\sum_{n=0}^{\infty}Q^{n}\left\langle\prod_{i=1}^{m}\tau_{k_{i}}(\mathsf{p_{1}})\prod_{j=m+1}^{r}\tau_{k_{j}+1}(1)\cdot\lambda_{g}\right\rangle_{g,n}^{\pi,\circ}=C\frac{(-1)^{g}}{24}\frac{|B_{2g}|}{4g}\left(Q\frac{d}{dQ}\right)^{m-1}E_{2g}(Q)\,,$
where
$C=\frac{(2g+r-3)!\sum_{i=1}^{m}(2k_{i}+1)}{2^{2g-2}(2g+m-2)!(2k_{1}+1)!!\ldots(2k_{r}+1)!!}\,.$
$\bullet$ For $\mathsf{F}=\lambda_{g-2}$ and $r=m=1$, we obtain (2.3):
$\langle\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}\rangle^{E,\circ}_{g,n}=(-1)^{g-1}\frac{4g}{B_{2g}}\sigma_{2g-1}(n)\int_{\overline{\mathcal{M}}_{g,1}}\psi_{1}\lambda_{g}\lambda_{g-1}\lambda_{g-2}=\frac{|B_{2g-2}|}{(2g-2)!}\sigma_{2g-1}(n)\,.$
Further development of these ideas will appear in [23].
3. The invariant $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$
3.1. Overview
We present here the proof of Theorem 1:
$\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=-\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\left(\mathsf{Tr}_{n}+\sum_{k=2}^{n-1}{\sigma_{-1}(n-k)}\mathsf{Tr}_{k}\right)\,.$
Our strategy is to convert the invariant $\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ to a families Hodge integral. After several steps related to the connected/disconnected calculus and descendent/relative correspondence, we will show that Theorem 1 follows from Theorem 2.
3.2. Connected/disconnected calculus
The genus 1 Gromov-Witten invariants of ${\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ are expressed in terms of families invariants with possibly disconnected [^20] The superscript $\bullet$ indicates possibly disconnected domain curves (but no connected component of the domain is contracted to a point). domain curves by by (0.3):
$\langle\,D\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=-\langle(2,1^{n-2})\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\frac{1}{i}\langle(2,1^{n-2})\rangle^{\pi_{\mathbb{C}^{2}},\bullet}\,.$
We will transform $\langle(2,1^{n-2})\rangle^{\pi_{\mathbb{C}^{2}},\bullet}$ to invariants with connected domain curves. The connected/disconnected correspondence is well-known for the Gromov-Witten theory of a fixed target. For families invariants, new aspects appear.
There are two connected Gromov-Witten invariants ((i) and (iii)) and two partition counts ((ii) and (iv)) which occur in the connected/disconnect calculus for $\langle(2,1^{n-2})\rangle^{\pi_{\mathbb{C}^{2}},\bullet}$:
(i) Let ¯$\overline{\mathcal{M}}^{\circ}_{g}(\pi,(2,1^{n-2}))$ be the moduli space of degree $n$ stable relative maps to the fibers of the universal elliptic curve
$\pi:\mathcal{E}\rightarrow\overline{\mathcal{M}}_{1,1}$
with connected [^21] The superscript $\circ$ denotes connected domains. domains of genus $g$ and relative condition at the marking given by $(2,1^{n-2})$. Let
$\langle\lambda_{g-2}\lambda_{g}|(2,1^{n-2})\rangle_{g,n}^{\pi,\circ}=\int_{[\overline{\mathcal{M}}^{\circ}_{g}(\pi,(2,1^{n-2}))]^{vir_{\pi}}}\lambda_{g-2}\lambda_{g}\,.$
The descendent/relative correspondence via the degeneration to the normal cone of the section $\mathsf{p}_{1}$ yields the following result proven in Section 3.5.
Proposition 8.
For $g\geq 2$ and $n\geq 2$, we have
$\langle\tau_{1}(\mathsf{p}_{1})\lambda_{g-2}\lambda_{g}\rangle_{g,n}^{\pi,\circ}=\frac{\sigma_{1}(n)}{24}\frac{|B_{2g-2}|}{(2g-2)!}+\langle\lambda_{g-2}\lambda_{g}|(2,1^{n-2})\rangle_{g,n}^{\pi,\circ}\,.$
Since $\langle\tau_{1}(\mathsf{p}_{1})\lambda_{g-2}\lambda_{g}\rangle_{g,n}^{\pi,\circ}$ was calculated in Theorem 2, Proposition 8 completely determines the invariant $\langle\lambda_{g-2}\lambda_{g}|(2,1^{n-2})\rangle_{g,n}^{\pi,\circ}$.
The $g=1$ case takes a special form. For $n\geq 2$, we easily obtain
$\langle\tau_{1}(\mathsf{p}_{1})\rangle_{1,n}^{\pi,\circ}=\frac{\sigma_{1}(n)}{24}\frac{|B_{0}|}{0!}+\langle(2,1^{n-2})\rangle_{1,n}^{\pi,\circ}\,.$
Using the evaluation (0.6) of $\langle\tau_{1}(\mathsf{p}_{1})\rangle_{1,n}^{\pi,\circ}$, we see
$\langle(2,1^{n-2})\rangle_{1,n}^{\pi,\circ}=0$
for $n\geq 2$.
(ii) For an integer $l\geq 1$, let $\mathsf{Part}(l)$ be the number of partitions of $l$. Partitions arise naturally in the Gromov-Witten theory of $E$: $\mathsf{Part}(l)$ is the count of possibly disconnected unramified covers of $E$ of degree $l$ where each cover is weighted by the order of the automorphism group. The corresponding generating series is
$\mathcal{P}(x)=1+\sum_{l=1}^{\infty}\mathsf{Part}(l)\,x^{l}=\prod_{l=1}^{\infty}\frac{1}{1-x^{l}}\,.$
(iii) Let $E$ be a fixed elliptic curve. We denote by $\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\circ}$ the connected genus $g$, degree $n$, $\mathsf{T}$-equivariant Gromov-Witten invariant of $E\times\mathbb{C}^{2}$ with relative condition at the divisor $\{p_{1}\}\times\mathbb{C}^{2}$ given by $(2,1^{n-2})$. The connected/disconnected correspondence here is
(3.1) $\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}=\sum_{2\leq m\leq n}\langle(2,1^{m-2})\rangle_{g,m}^{E\times\mathbb{C}^{2},\circ}\cdot\mathsf{Part}(n-m)\,.$
The disconnected invariants $\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}$ can be calculated by degenerating $E$ to a curve of arithmetic genus $1$ with a unique node and applying the correspondence between local Gromov-Witten theory of $\mathbb{P}^{1}$ and quantum cohomology ring of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ [3, 37]:
(3.2) $-\sum_{g\in\mathbb{Z}}u^{2g-3}\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}=(-i)\cdot\mathsf{trace}\left(\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)\right)=(-i)\cdot\mathsf{Tr}_{n}\cdot(t_{1}+t_{2}),$
after $-q=e^{iu}$.
(iv) For an integer $l\geq 1$, let $\widetilde{\mathsf{Part}}(l)$ be the count of possibly disconnected unramified covers of $E$ of degree $l$ where each cover is weighted by the order of automorphism group and the number of connected components. For example, $\widetilde{\mathsf{Part}}(1)=1$ and
$\widetilde{\mathsf{Part}}(2)=1+3/2=5/2\,,$
where $1=(1/2)\cdot 1\cdot 2$ is the contribution of the disconnected cover and $3/2=(1/2)\cdot 3\cdot 1$ is the contribution of the connected covers. The generating series is
$\widetilde{\mathcal{P}}(x)=\sum_{l=1}^{\infty}\widetilde{\mathsf{Part}}(l)\,x^{l}\,.$
Lemma 9.
The series $\widetilde{\mathcal{P}}$ is determined by the equation $\widetilde{\mathcal{P}}=\mathcal{P}\operatorname{log}\mathcal{P}$.
Proof.
Let $\mathsf{Hur}(l,k)$ be the automorphism-weighted count of possibly disconnected unramified covers of an elliptic curve of degree $l$ with exactly $k$ connected component, and let
$\mathcal{F}(x,y)=1+\sum_{l=1}^{\infty}\sum_{k=1}^{\infty}\operatorname{\mathsf{Hur}}(l,k)\,x^{l}y^{k}\,.$
Since $\operatorname{log}(\mathcal{P}(x))$ is the generating series of connected unramified covers of an elliptic curve,
$\mathcal{F}(x,y)=\operatorname{exp}(y\operatorname{log}(\mathcal{P}(x)))\,.$
We then obtain $\widetilde{\mathcal{P}}(x)=\partial_{y}\mathcal{F}(x,y)|_{y=1}=\mathcal{P}(x)\operatorname{log}\mathcal{P}(x)$ by the definition of $\mathsf{Hur}(l,k)$. ∎
We can now state the main connected/disconnect equation which will play an essential role in the proof of Theorem 1. The proof will be presented in Section 3.4.
Proposition 10 (Connected/disconnected calculus).
For $g\geq 2$ and $n\geq 2$, we have
$\langle(2,1^{n-2})\rangle_{g,n}^{\pi_{\mathbb{C}^{2}},\bullet}$ $=$ $\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\sum_{2\leq m\leq n}\langle\lambda_{g-2}\lambda_{g}|(2,1^{m-2})\rangle_{g,m}^{\pi,\circ}\cdot\mathsf{Part}(n-m)$
$-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}\sum_{2\leq m\leq n}\langle(2,1^{m-2})\rangle_{g,m}^{E\times\mathbb{C}^{2},\circ}\cdot\widetilde{\mathsf{Part}}(n-m)\,.$
The $g=1$ case of the connected/disconnected calculus takes a special form. For $n\geq 2$,
(3.3) $\langle(2,1^{n-2})\rangle_{1,n}^{\pi_{\mathbb{C}^{2}},\bullet}=-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}\sum_{2\leq m\leq n}\langle(2,1^{m-2})\rangle_{1,m}^{E\times\mathbb{C}^{2},\circ}\cdot\widetilde{\mathsf{Part}}(n-m)\,.$
The genus $g=1$ case will be discussed in Section 3.4. In fact, both sides of (3.3) vanish for $g=1$.
3.3. Proof of Theorem 1
Let $\mathcal{B}(u,Q)$ denote the power series
$\mathcal{B}(u,Q)=\sum_{g=1}^{\infty}\sum_{m=1}^{\infty}\frac{|B_{2g-2}|}{(2g-2)!}(\sigma_{2g-1}(m)-\sigma_{1}(m))Q^{m}u^{2g-3}.$
The $u^{-1}$ term of $\mathcal{B}(u,y)$ corresponding to $g=1$ vanishes. Recall the definition
$\mathsf{Tr}_{m}=\frac{1}{t_{1}+t_{2}}\mathsf{trace}(\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})})\,$
of Section 0.4.1.
Lemma 11.
Under the variable change $-q=e^{iu}$, we have
$(-i)\sum_{m=1}^{\infty}\mathsf{Tr}_{m}(q)\,Q^{m}=\mathcal{P}(Q)\mathcal{B}(u,Q).$
Proof.
The diagonal terms of $\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}$ have been described in [37, Section 2]:
$\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}=\sum_{r=1}^{\infty}\left(\frac{r}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}-\frac{1}{2}\frac{(-q)+1}{(-q)-1}\right)\alpha_{-r}\alpha_{r}+\ldots,$
where $\alpha_{-r}$ and $\alpha_{r}$ are the standard creation and annihilation operators on the subspace $\mathcal{F}_{m}\subset\mathcal{F}$ of Fock space
$\mathcal{F}=\sum_{m=0}^{\infty}\mathcal{F}_{m}\,.$
Using the diagonal elements, we compute
(3.4) $(-i)\sum_{m=1}^{\infty}\mathsf{Tr}_{m}(q)\,Q^{m}=\sum_{r\geq 1}\left(\frac{-ir}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}-\frac{-i}{2}\frac{(-q)+1}{(-q)-1}\right)\sum_{m=1}^{\infty}\mathsf{trace}(\alpha_{-r}\alpha_{r}\,{}_{\mid\mathcal{F}_{m}})\cdot Q^{m}\,.$
Since $\alpha_{-r}\alpha_{r}(|\mu\rangle)=r\cdot|\{i\mid\mu_{i}=r\}|\cdot|\mu\rangle$,
(3.5) $\sum_{m=1}^{\infty}\mathsf{trace}(\alpha_{-r}\alpha_{r}\,{}_{\mid\mathcal{F}_{m}})\cdot Q^{m}=r\frac{Q^{r}}{1-Q^{r}}\prod_{m=1}^{\infty}\frac{1}{1-Q^{m}}\,.$
The expansion of the cotangent function yields
(3.6) $\begin{split}\frac{-ir}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}&=\frac{-ir}{2}\frac{e^{iur/2}+e^{-iur/2}}{e^{iur/2}-e^{iur/2}}\\ &=-\frac{r}{2}\cot\left(\frac{ur}{2}\right)\\ &=-\frac{1}{u}+\sum_{h=0}^{\infty}\frac{|B_{2h}|r^{2h}}{(2h)!}u^{2h-1}\,.\end{split}$
After putting together (3.4), (3.5) and (3.6) and shifting $h=g-1$, we obtain:
$\begin{split}(-i)\sum_{m=1}^{\infty}\mathsf{Tr}_{m}(q)\,Q^{m}&=\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\sum_{r=1}^{\infty}\frac{|B_{2g-2}|}{(2g-2)!}(r^{2g-1}-r)\frac{Q^{r}}{1-Q^{r}}u^{2g-3}\right)\\ &=\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\frac{|B_{2g-2}|}{(2g-2)!}u^{2g-3}\sum_{r=1}^{\infty}(r^{2g-1}-r)(Q^{r}+Q^{2r}+\ldots)\right)\\ &=\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\frac{|B_{2g-2}|}{(2g-2)!}u^{2g-3}\sum_{m=1}^{\infty}Q^{m}\sum_{k\mid m}(k^{2g-1}-k)\right)\\ &=\mathcal{P}(Q)\mathcal{B}(u,Q)\,.\end{split}$
By Theorem 2 and Proposition 8, we obtain
(3.7) $\sum_{g=1}^{\infty}\sum_{n=1}^{\infty}u^{2g-3}Q^{n}\langle\lambda_{g-2}\lambda_{g}|(2,1^{n-2})\rangle_{g,n}^{\pi,\circ}=\frac{1}{24}\mathcal{B}(u,Q).$
The $g=1$ terms on the left are degenerate and interpreted as
$\langle(2,1^{n-2})\rangle_{1,n}^{\pi,\circ}=0\,,$
so all $u^{-1}$ vanish on both sides of (3.7).
By Lemma 11 and (3.2), the connected/disconnected equation for a fixed elliptic curve target (3.1) can be written as
$\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\sum_{n=1}^{\infty}\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\circ}u^{2g-3}Q^{n}\right)=i(t_{1}+t_{2})\sum_{n=1}^{\infty}\mathsf{Tr}_{n}(q)\,Q^{n}=-(t_{1}+t_{2})\mathcal{P}(Q)\mathcal{B}(u,Q)\ ,$
after $-q=e^{iu}$. The connected/disconnected calculus of Proposition 10 then yields
$\begin{split}\sum_{g=1}^{\infty}\sum_{n=1}^{\infty}\langle(2,1^{n-2})\rangle_{g,n}^{\pi_{\mathbb{C}^{2}},\bullet}Q^{n}u^{2g-3}&=-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}(-(t_{1}+t_{2})\mathcal{B}(u,Q))\cdot\widetilde{\mathcal{P}}(Q)+\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\frac{1}{24}\mathcal{B}(u,Q)\mathcal{P}(Q)\\ &=\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\mathcal{B}(u,Q)\mathcal{P}(Q)(1+\log\mathcal{P}(Q))\\ &=\frac{1}{i}\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\left(\sum_{n=1}^{\infty}\mathsf{Tr}_{n}(q)\,Q^{n}\right)\left(1+\sum_{k=1}^{\infty}\sigma_{-1}(k)Q^{k}\right)\,.\end{split}\,$
In the second equality, we have applied Lemma 9. In the third equality, we have used
$\operatorname{log}\mathcal{P}(Q)=-\sum_{n=1}^{\infty}\operatorname{log}(1-Q^{n})=\sum_{k=1}^{\infty}\sigma_{-1}(k)Q^{k}\,.$
Since, by Theorem A,
$\langle D\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\frac{1}{i}\langle(2,1^{n-2})\rangle^{\pi_{\mathbb{C}^{2}},\bullet}\,,$
$\langle D\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is $\frac{-1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}$ times the $Q^{n}$ coefficient of
$\left(\sum_{n=1}^{\infty}\mathsf{Tr}_{n}(q)\,Q^{n}\right)\left(1+\sum_{k=1}^{\infty}\sigma_{-1}(k)Q^{k}\right)\,.$
We have proven Theorem 1. ∎
3.4. Proof of Proposition 10
We study here the connected/disconnected calculus for the family $\pi_{\mathbb{C}^{2}}$. To start, we consider the universal elliptic curve
$\pi:\mathcal{E}\rightarrow\overline{\mathcal{M}}_{1,1}\,$
with section $\mathsf{p}_{1}$. Let $(m,\mathbf{k},\mathbf{g})$ be a triple,
  • $2\leq m\leq n$,
  • $\mathbf{k}=(k_{1},\ldots,k_{s})$ is a partition of $n-m$,
  • $\mathbf{g}=(g_{0},\ldots,g_{s})$ is a partition of $g+s-1$,
where the parts of $\mathbf{k}$ and $\mathbf{g}$ are positive integers. We define
¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ =¯$=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{1}}^{\circ}(\pi,(1^{k_{1}}))\times_{\overline{\mathcal{M}}_{1,1}}\ldots\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}})),$
¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ =¯$=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times\overline{\mathcal{M}}^{\circ}_{g_{1}}(\pi,(1^{k_{1}}))\times\ldots\times\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}}))\,,$
where ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ is the base change of ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ along the diagonal $\Delta:\overline{\mathcal{M}}_{1,1}\to\overline{\mathcal{M}}_{1,1}\times\ldots\times\overline{\mathcal{M}}_{1,1}$.
The fixed locus of ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},n,(2,1^{n-2}))$ for the action of the torus $(\mathbb{C}^{*})^{2}$ is the disjoint union of the stacks ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$, up to an automorphism factor which we denote by $\operatorname{Aut}(\mathbf{k},\mathbf{g})$, and the restriction of the virtual class [¯$[\overline{\mathcal{M}}_{g}^{\bullet}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}$ to ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ agrees with $\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}$. By the localization formula of [18], we have
[¯$[\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\frac{\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}}{e_{(\mathbb{C}^{*})^{2}}(N^{vir})}\,,$
where the contribution of the virtual normal bundle on each factor of ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ is given by
(3.8) $\begin{split}\frac{1}{\mathsf{e}_{(\mathbb{C}^{*})^{2}}(N^{vir})}&=\frac{c(\mathbb{E}^{\vee}\otimes t_{1})c(\mathbb{E}^{\vee}\otimes t_{2})}{t_{1}t_{2}}\\ &=-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-1}+\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-2}-(t_{1}+t_{2})\lambda_{g_{i}-1}\lambda_{g_{i}-2}+\ldots\end{split}$
if $g_{i}>1$, and $-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{1}+1$ if $g_{i}=1$. We denote the class (3.8) by $N_{g_{i}}$, and note that the lower order terms in the Hodge classes will not contribute for dimensional reasons. After integration, we obtain
(3.9) $\begin{split}\langle(2,1^{n-2})\rangle_{g,n}^{\pi_{\mathbb{C}^{2}},\bullet}&=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\int_{[\overline{\mathcal{M}}^{\mathsf{fiber}}_{(m,\mathbf{k},\mathbf{g})}]^{vir}}\prod_{i=0}^{s}N_{g_{i}}\\ &=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\sum_{j=0}^{s}\int_{[\overline{\mathcal{M}}_{(m,\mathbf{k},\mathbf{g})}]^{vir}}\prod_{i=0}^{s}N_{g_{i}}\prod_{i\neq j}\operatorname{ev}_{i}^{*}([\mathsf{pt}])\\ &=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\left(\langle N_{g_{0}}|(2,1^{m-2})\rangle_{g_{0},m}^{\pi,\circ}\prod_{i=1}^{s}\langle N_{g_{i}}|(1^{k_{i}})\rangle_{g_{i},k_{i}}^{E,\circ}\right.\\ &+\left.\langle N_{g_{0}}|(2,1^{m-2})\rangle_{g_{0},m}^{E,\circ}\sum_{j=1}^{s}-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\langle\lambda_{g_{j}}\lambda_{g_{j}-1}|(1^{k_{j}})\rangle_{g_{j},k_{j}}^{\pi,\circ}\prod_{i\neq j}\langle N_{g_{i}}|(1)\rangle_{g_{i},k_{i}}^{E,\circ}\right)\end{split}$
where $\operatorname{ev}_{i}:\overline{\mathcal{M}}_{(m,\mathbf{k},\mathbf{g})}\to\overline{\mathcal{M}}_{1,1}$ recovers the isomorphism class of the target curve of the $i$-th map. The second equality follows from the decomposition of the diagonal of ¯$\overline{\mathcal{M}}_{1,1}$:
$[\Delta]=\sum_{j=0}^{s}[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\boxtimes 1\boxtimes[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\in\operatorname{\mathsf{CH}}^{s}(\overline{\mathcal{M}}_{1,1}\times\cdots\times\overline{\mathcal{M}}_{1,1})\,.$
If $g_{i}>1$, the class $N_{g_{i}}$ contains a $\lambda_{g_{i}}$ factor. Hence, $\langle N_{g_{i}}|(1^{k_{i}})\rangle_{g_{i},k_{i}}^{E,\circ}=0$ unless $g_{i}=1$, in which case
$\langle N_{1}|(1^{k_{i}})\rangle_{1,k_{i}}^{E,\circ}=\langle(1^{k_{i}})\rangle_{1,k_{i}}^{E,\circ}=\frac{\sigma(k_{i})}{k_{i}}\,.$
Lemma 12.
If $g_{i}>1$, we have $\langle\lambda_{g_{i}}\lambda_{g_{i}-1}|(1^{k_{i}})\rangle_{g_{i},k_{i}}^{\pi,\circ}=0$. If $g_{i}=1$,
$\langle\lambda_{1}\lambda_{0}|(1^{k_{i}})\rangle_{1,k_{i}}^{\pi,\circ}=\frac{1}{24}\langle(1^{k_{i}})\rangle_{1,k_{i}}^{E,\circ}=\frac{\sigma_{1}(k_{i})}{24k_{i}}\,.$
Proof.
The first step of the argument is the equality
$\langle\lambda_{g_{i}}\lambda_{g_{i}-1}|(1^{k_{i}})\rangle_{g_{i},k_{i}}^{\pi,\circ}=\langle\lambda_{g_{i}}\lambda_{g_{i}-1}\rangle_{g_{i},k_{i}}^{\pi,\circ}$
obtained from the standard degeneration to the normal cone of the section $\mathsf{p}_{1}$.
We now follow the geometric notation and analysis of Section 2.5. The moduli space of stable maps to the fibers of $\pi$ lies over ¯$\overline{\mathcal{M}}_{1,1}$,
$\epsilon:\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\rightarrow\overline{\mathcal{M}}_{1,1}\,.$
The universal curve $\mu:\mathcal{C}\rightarrow\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$ carries a universal evaluation map
$\mathcal{C}$$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$$f$$\mu$$\nu$
to the universally expanded target
$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\times_{\overline{\mathcal{M}}_{1,1}}\mathcal{E}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\,.$$h$$\nu$$\pi_{M}$
[…]
(3.10) $\epsilon^{*}\mathbb{E}_{1}\cong\nu_{*}\omega_{\mu}\ \stackrel{{\scriptstyle f^{*}}}{{\rightarrow}}\ \mu_{*}\omega_{\mu}\cong\mathbb{E}_{g}$
over ¯$\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$. Since $k_{i}>0$, the map (3.10) is injective, so we obtain an exact sequence
$0\rightarrow\epsilon^{*}\mathbb{E}_{1}\rightarrow\mathbb{E}_{g}\rightarrow\mathbb{F}\rightarrow 0\,,$
where $\mathbb{F}$ is a rank $g-1$ vector bundle on ¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$. We therefore have a factorization
(3.11) $\lambda_{g}\ =\ c_{1}(\epsilon^{*}\mathbb{E}_{1})\cdot c_{g-1}(\mathbb{F})\ =\ \frac{1}{24}\epsilon^{*}([E])\cdot\lambda_{g-1}\,$
on ¯$\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$, where $[E]\in\overline{\mathcal{M}}_{1,1}$ is the moduli point of a fixed nonsingular elliptic curve $E$.
The factorization (3.11) of $\lambda_{g}$ implies
(3.12) [¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})]^{vir}}\lambda_{g_{i}}\lambda_{g_{i}-1}=\frac{1}{24}\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(E,k_{i})]^{vir}}\lambda^{2}_{g_{i}-1}\,.$
If $g_{i}>1$, then $\lambda_{g_{i}-1}^{2}=2\lambda_{g_{i}}\lambda_{g_{i}-2}$. Hence, the integral (3.12) is 0 by $\lambda_{g}$-vanishing for elliptic targets. If $g_{i}=1$, the integral is easily evaluated as claimed. ∎
Therefore, the only positive contributions to the right hand side of (3.9) come from triplets $(e,\mathbf{k},\mathbf{g})$ where $\mathbf{g}=(g,1,\ldots,1)$.
$\bullet$ For $g>1$, we have
$\langle N_{g}|(2,1^{m-2})\rangle_{g,m}^{\pi,\circ}=\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\langle\lambda_{g}\lambda_{g-2}|(2,1^{m-2})\rangle_{g,m}^{\pi,\circ}\,.$
By a similar localization analysis, we have
$\langle N_{g}|(2,1^{m-2})\rangle_{g,m}^{E,\circ}=-(t_{1}+t_{2})\langle\lambda_{g-1}\lambda_{g-2}|(2,1^{m-2})\rangle_{g,m}^{E,\circ}=\langle(2,1^{m-2})\rangle^{E\times\mathbb{C}^{2}}_{g,m}$
Therefore, the formula in (3.9) simplifies to
$\begin{split}\langle(2,1^{n-2})\rangle_{g,n}^{\pi_{\mathbb{C}^{2}},\bullet}=\sum_{m=2}^{n}\sum_{\mathbf{k}\,\vdash n-m}&\left(\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\langle\lambda_{g}\lambda_{g-2}|(2,1^{m-2})\rangle_{g,m}^{\pi,\circ}\frac{\prod_{i}\langle(1^{k_{i}})\rangle_{1,k_{i}}^{E,\circ}}{\operatorname{Aut}(\mathbf{k})}\right.\\ &\phantom{+}\left.-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}\langle(2,1^{m-2})\rangle_{g,m}^{E\times\mathbb{C}^{2},\circ}\frac{l(\mathbf{k})\prod_{i}\langle(1^{k_{i}})\rangle_{1,k_{i}}^{E,\circ}}{\operatorname{Aut}(\mathbf{k})}\right).\end{split}$
Summing over all possible partitions, we obtain the coefficients $\operatorname{\mathsf{Part}}(n-m)$ and $\widetilde{\operatorname{\mathsf{Part}}}(n-m)$ of Proposition 10 as defined in Section 3.2.
$\bullet$ For $g=1$, $\langle N_{1}|(2,1^{m-2})\rangle_{1,m}^{\pi,\circ}=\langle(2,1^{m-2})\rangle_{1,m}^{\pi,\circ}=0$ by Proposition 8, so both sides of formula (3.3) after Proposition 10 vanish. ∎
3.5. Proof of Proposition 8
Let $g\geq 2$. Let $\operatorname{Bl_{\mathsf{p}_{1}\times\{0\}}(\mathcal{E}\times\mathbb{A}^{1})}$ be the degeneration to the normal cone of the section $\mathsf{p}_{1}$. The special fiber over $0\in\mathbb{A}^{1}$ is the family
$\pi\cup\pi_{P}:\mathcal{E}\cup\mathbb{P}(T_{\mathsf{p_{1}}}\mathcal{E}\oplus\mathbb{C})\longrightarrow\overline{\mathcal{M}}_{1,1},$
where the gluing identifies the section $\mathsf{p}_{1}$ of $\mathcal{E}$ with the 00-section of $\mathbb{P}(T_{\mathsf{p}_{1}}\mathcal{E}\oplus\mathbb{C})$. The universal section $\mathsf{p}_{1}$ is now the $\infty$-section of $\pi_{P}$. We will use the degeneration formula
(3.13) $\langle\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\rangle_{g,n}^{\pi,\circ}=\sum_{(\Gamma_{1},\Gamma_{2},\mu)}\langle\Lambda_{1}|\mu\rangle_{\Gamma_{1}}^{\pi,\bullet}\langle\Lambda_{2}\tau_{1}(\infty)|\mu^{*}\rangle_{\Gamma_{2}}^{\pi_{P},\bullet}\,,$
with notation:
  1. $\bullet$
    $\Gamma_{1}$ and $\Gamma_{2}$ are possibly disconnected topological types corresponding to the domains mapping to the components $\mathcal{E}$ and $\mathbb{P}(T_{\mathsf{p}}\mathcal{E}\oplus\mathbb{C})$,
  2. $\bullet$
    $\mu$ is a partition of $d$ decorated with elements of a basis of $H^{*}(\mathsf{p})=H^{*}(\overline{\mathcal{M}}_{1,1})$,
  3. $\bullet$
    $\Lambda_{1}$ and $\Lambda_{2}$ reflect the distribution of the Hodge insertions.
The Hodge insertion $\lambda_{g}\lambda_{g-2}=2\lambda_{g-1}^{2}$ annihilates all contributions of graphs which are not of compact type or have more than two vertices with positive genus. If there are exactly two vertices of genera $0<g_{1},g_{2}<g$, then the Hodge insertion distributes as
(3.14) $\lambda_{g}\lambda_{g-2}|_{\overline{\mathcal{M}}_{g_{1}}\times\overline{\mathcal{M}}_{g_{2}}}=\lambda_{g_{1}}\lambda_{g_{1}-1}\boxtimes\lambda_{g_{2}}\lambda_{g_{2}-1}.$
An analysis of the virtual dimension of each vertex shows that all vertex contributions vanish for dimension reasons except for the following list:
(3.15) $\begin{split}\langle\lambda_{g}\lambda_{g-2}|(2[\overline{\mathcal{M}}_{1,1}],1[\overline{\mathcal{M}}_{1,1}],\ldots,1[\overline{\mathcal{M}}_{1,1}])\rangle_{g}^{\pi,\circ}\end{split}$
(3.16) $\begin{split}\langle\lambda_{g}\lambda_{g-2}|(1[\mathsf{pt}],1[\overline{\mathcal{M}}_{1,1}],\ldots,1[\overline{\mathcal{M}}_{1,1}])\rangle_{g}^{\pi,\circ}\end{split}$
(3.17) $\begin{split}\langle\lambda_{g_{1}}\lambda_{g_{1}-1}|(1[\overline{\mathcal{M}}_{1,1}],\ldots,1[\overline{\mathcal{M}}_{1,1}])\rangle_{g_{1}}^{\pi,\circ}\end{split}$
(3.18) $\begin{split}\langle(1[\mathsf{pt}])\rangle_{0}^{\pi_{P},\circ}\end{split}$
(3.19) $\begin{split}\langle\tau_{1}(\infty)|(1[\overline{\mathcal{M}}_{1,1}])\rangle_{0}^{\pi_{P},\circ}\end{split}$
(3.20) $\begin{split}\langle\tau_{1}(\infty)|(2[\mathsf{pt}])\rangle_{0}^{\pi_{P},\circ}\end{split}$
(3.21) $\begin{split}\langle\tau_{1}(\infty)|(1[\mathsf{pt}],1[\mathsf{pt}])\rangle_{0}^{\pi_{P},\circ}\end{split}$
(3.22) $\begin{split}\langle\tau_{1}(\infty)\lambda_{g_{2}}\lambda_{g_{2}-1}|(1[\mathsf{pt}])\rangle_{g_{2}}^{\pi_{P},\circ}\end{split}$
Three further vanishings hold for other reasons:
  1. $\bullet$
    Vertex (3.16) vanishes because
    [¯$[\overline{\mathcal{M}}_{g}(\pi/\mathsf{p}_{1})]^{vir}\cap\operatorname{ev}_{\overline{\mathcal{M}}_{1,1}}^{*}(\mathsf{pt})=[\overline{\mathcal{M}}_{g}(E/p_{1})]^{vir}$
    pairs to zero with $\lambda_{g}$.
  2. $\bullet$
    Vertex (3.21) vanishes because of the two point conditions over ¯$\overline{\mathcal{M}}_{1,1}$.
  3. $\bullet$
    Vertex (3.17) vanishes unless $g_{1}=1$ by Lemma 12.
It follows that the only possible combinatorial types that contribute to the right hand side of (3.13) are given by the two configurations of Figure 1, where the first configuration is counted $n$ times.
[…]
Figure 1. Non-zero contributions to (3.13)
The vertices (3.18) and (3.20) contribution factors $1$ and $1/2$ respectively (the latter is cancelled by the multiplicity which occurs in the degeneration formula).
The above vertex analysis show that (3.13) specializes to:
$\langle\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\rangle_{g,n}^{\pi,\circ}=\frac{\sigma_{1}(n)}{24}\langle\tau_{1}(\infty)\lambda_{g-1}\lambda_{g-2}|(1[\mathsf{pt}])\rangle_{g-1,1}^{\pi_{P}}+\langle\lambda_{g}\lambda_{g-2}|(2,1^{n-2})\rangle_{g,n}^{\pi,\circ}\,.$
To complete the proof of Proposition 8, we must evaluate
$\langle\tau_{1}(\infty)\lambda_{g-1}\lambda_{g-2}|(1[\mathsf{pt}])\rangle_{g-1,1}^{\pi_{P}}=\langle\tau_{1}(\infty)\lambda_{g-1}\lambda_{g-2}\rangle_{g-1,1}^{\mathbb{P}^{1}}\,,$
where the equality with the absolute invariant is proven by the standard degeneration method. To calculate, we localize with respect to the $\mathbb{C}^{*}$- action on $\mathbb{P}^{1}$. Since the integrand has a $\lambda_{g}\lambda_{g-1}$ insertion, there are only two components of the fixed locus that contribute, corresponding to the graphs in Figure 2, see [18].
[…]
Figure 2. Localization contributions, where half edges correspond to marked points, vertices are contracted components, and edges are non-contracted components.
After expanding the localization formula, we conclude
$\langle\tau_{1}(\infty)\lambda_{g-1}\lambda_{g-2}\rangle_{g-1,1}^{\mathbb{P}^{1}}=(2g-2)\int_{\overline{\mathcal{M}}_{g-1,1}}\frac{\lambda_{g-1}\lambda_{g-2}c(\mathbb{E}^{\vee})}{1-\psi}=\frac{|B_{2g-2}|}{(2g-2)!}\,,$
where the Hodge integral on the right is calculated in [41]. ∎
4. Reconstruction of multi-point invariants
4.1. Reduction to $1$-point series
We recall the notation introduced in Section 0.1:
(4.1) $\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\sum_{d=0}^{\infty}\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\rangle_{1,d}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}q^{d}\,.$
In order to emphasize the number of insertions, we also write
$\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,{\framebox{{\tiny$\ell$}}}}\,,$
where the boxed subscript indicates insertion number (not the curve degree as in (4.1)).
We prove here the reduction of the genus $1$ Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to $1$-point series as stated in Theorem 4:
To every genus $1$ series $\big{\langle}D^{\star k_{1}},...,D^{\star k_{\ell}}\big{\rangle}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$, there are canonically associated functions
[…]
for which the following equation holds:
$\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\sum_{k=0}^{|\mathsf{Part}(n)|-1}\,\sum_{m=0}^{\ell-1}\,C_{k,m}\cdot\left(q\frac{d}{dq}\right)^{m}\big{\langle}D^{*k}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\,.$
4.2. Proof of Theorem 4
We adapt the reconstruction strategy of [50, Section 6] to prove the result. Let $m\geq 0$, and fix classes
$v_{1},v_{2},v_{3},v_{4},\phi_{1},...,\phi_{m}\in H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))\,.$
Pulling back Getzler’s relation from ¯$\overline{\mathcal{M}}_{1,4}$ to ¯$\overline{\mathcal{M}}_{1,4+m}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),d)$ and summing over the curve degree $d$, we obtain the following equation (see [26, Equation (27)])
[…]
where the relation $\mathsf{X}\sim\mathsf{Y}$ means that the difference $\mathsf{X}-\mathsf{Y}$ is an explicit sum of arbitrary genus 00 invariants and genus $1$ invariants with at most $m+1$ insertions. Getzler’s relation can be equivalently cast in the following form:
$\Psi(v_{1},v_{2},v_{3},v_{4})$ $=$ $\ \ \,\langle v_{1}\star v_{2},v_{3}\star v_{4},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}+\langle v_{1}\star v_{3},v_{2}\star v_{4},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$
$+\langle v_{1}\star v_{4},v_{2}\star v_{4},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$
$-\langle v_{1},v_{2}\star v_{3}\star v_{4},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}-\langle v_{2},v_{1}\star v_{3}\star v_{4},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$
$-\langle v_{3},v_{1}\star v_{2}\star v_{4},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}-\langle v_{4},v_{1}\star v_{2}\star v_{3},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$
satisfies the relation
(4.2) $\Psi(v_{1},v_{2},v_{3},v_{4})\sim 0\,.$
Let $a,b\in H^{\leq 2}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$, and let $\phi_{1},...,\phi_{m}\in H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$. For integers $l\geq 2$ and $0\leq i\leq l$ define the following series:
$h(l)$ $=$ $\langle a\star b\star D^{*(l-2)},D^{*2},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\,,$
$f(i)$ $=$ $\langle a\star D^{*i},b\star D^{*(l-i)},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\,.$
From the string and divisor equations, we immediately obtain
$f(0)\sim 0\,,\quad f(l)\sim 0\,.$
By expanding the definitions, we see, for $20\leq k\leq l-$ 2,
$-\Psi(a\star D^{*k},b\star D^{*(l-2-k)},D,D)$ $=$ $f(k+2)-2f(k+1)+f(k)-h(l)$
$\ \ \ +2\langle D,a\star b\star D^{*(l-1)},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\,.$
By applying (4.2) and the divisor equation, we obtain a difference equation,
$f(k+2)-2f(k+1)+f(k)-h(l)\sim 0\,,$
from (4.2). A linear algebraic result [50, Lemma 6.2] about solutions to the difference equation then yields
(4.4) $f(i)\sim-\frac{i(l-i)}{2}h(l)\,.$
for $0\leq i\leq l$.
We can write the relation (4.4) in the following form. Let $l\geq 2$, and let $i+j=l$ for $i,j\geq 0$. Then,
(4.5) $f(i)\sim-\frac{ij}{2}h(l)\,.$
The case $j=2$ of (4.5) with $b=1$ is
$\langle a\star D^{*i},D^{*2},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\sim-i\langle a\star D^{*i},D^{*2},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}.$
The case $j=1$ of (4.5) with $b=D$ is
$\langle a\star D^{*i},D\star D^{*1},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\sim-\frac{i}{2}\langle a\star D\star D^{*(i-1)},D^{*2},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}.$
These two relations imply
(4.6) $\langle D^{*i},D^{*2},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\sim 0$
for all $i\geq 1$. The $i=0$ case, $\langle D^{*0},D^{*2},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\sim 0$ follows from the string equation.
As a consequence of (4.6), we have $h(l)\sim 0$ for all $l\geq 2$. Then, relation (4.5) yields
$\langle a\star D^{*i},b\star D^{*j},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\sim 0\,$
for all $i,j\geq 0$ with $i+j\geq 2$. Again the $i+j=0$ and $i+j=1$ cases follow from the string equation. After setting $a=b=1$, we conclude
(4.7) $\langle D^{*i},D^{*j},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\sim 0\,$
for all $i,j\geq 0$.
Repeated use of relation (4.7) shows that $\langle D^{*i},D^{*j},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$ can be expressed in terms of genus 00 invariants and $1$-point genus $1$-functions. Moreover, since we use the divisor equation in the argument, the resulting expressions are linear in $1$-point genus $1$ functions and their $q\frac{d}{dq}$ derivatives. By the genus 00 reconstruction result in [37], the genus 00 series can be written as rational functions in matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$. The proof of Theorem 4 is complete. ∎
4.3. Givental’s formula in genus 1
Consider the (full) genus $1$ Gromov-Witten potential
$\mathcal{F}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(\mathsf{t},q)=\sum_{d\geq 0}\sum_{\ell\geq 0}\frac{q^{d}}{\ell!}\langle\mathsf{t},...,\mathsf{t}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{{\tiny$\ell$}}\,,d},\quad\mathsf{t}\in H_{\mathsf{T}}^{*}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))\,.$
We have, for $v_{1},...,v_{k}\in H_{\mathsf{T}}^{*}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$,
$\langle v_{1},...,v_{k}\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\nabla_{v_{1}}\nabla_{v_{2}}...\nabla_{v_{k}}\mathcal{F}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(\mathsf{t},q)\big{|}_{\mathsf{t}=0}\,.$
Since the quantum cohomology of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is semisimple, we can apply Givental’s formula [14]:
$\nabla_{v}\mathcal{F}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(\mathsf{t},q)=\frac{1}{2}\sum_{i=1}^{|\mathsf{Part}(n)|}\mathsf{R}^{1}_{ii}\cdot\nabla_{v}u_{i}+\frac{1}{48}\nabla_{v}\log\Big{(}\,\prod_{i=1}^{|\mathsf{Part(n)}|}\Delta_{i}\Big{)}\,.$
We follow here the notation of [14] and refer the reader to [14, 24] for an exposition:
  1. $\bullet$
    $\mathsf{R}^{1}_{ii}$ are matrix coefficients of the first term of the classifying $\mathsf{R}$-matrix,
    $\mathsf{R}=\mathsf{Id}+\mathsf{R}^{1}\cdot z+\mathsf{R}^{2}\cdot z^{2}+...\,,$
  2. $\bullet$
    $u_{1},\ldots,u_{|\mathsf{Part}(n)|}$ are the canonical coordinates,
  3. $\bullet$
    $\Delta_{1},\ldots,\Delta_{|\mathsf{Part}(n)|}$ are the inverses of the squares of the lengths of the corresponding idempotents $\epsilon_{1},\ldots,\epsilon_{|\mathsf{Part}(n)|}$.
By Theorem 4, we need only consider the 1-point invariants $\langle D^{\star k}\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$. Givental’s formula then can be written as:
(4.8) $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}=\frac{1}{2}\sum_{i=1}^{|\mathsf{Part}(n)|}\mathsf{R}^{1}_{ii}\,\big{|}_{\mathsf{t}=0}\cdot\nabla_{D^{\star k}}u_{i}\,\big{|}_{\mathsf{t}=0}+\frac{1}{48}\nabla_{D^{\star k}}\log\Big{(}\,\prod_{i=1}^{|\mathsf{Part(n)|}}\Delta_{i}\Big{)}\,\big{|}_{\mathsf{t}=0}\,.$
We explain first how to explicitly compute the functions $\Delta_{i}$. Denote by $\{e_{1},...,e_{|\mathsf{Part}(n)|}\}$ the distinct eigenvalues [37] of $D\star_{\mathsf{t}}$ and let $\{v_{1},...,v_{\mathsf{Part}(n)}\}$ be the corresponding eigenvectors. [^23] Here, the symbol $\star_{\mathsf{t}}$ denotes the big quantum product. We have
$v_{i}\star_{\mathsf{t}}v_{j}=\delta_{ij}c_{i}v_{i}\,,$
so the idempotents are $\epsilon_{i}=v_{i}/c_{i}$. By the Frobenius property,
$c_{i}\langle v_{i},1\rangle=\langle v_{i}\star_{\mathsf{t}}v_{i},1\rangle=\langle v_{i},v_{i}\rangle\,.$
We then compute:
$\Delta_{i}=\frac{1}{\langle\epsilon_{i},\epsilon_{i}\rangle}=\frac{c_{i}^{2}}{\langle v_{i},v_{i}\rangle}=\frac{1}{\langle v_{i},v_{i}\rangle}\cdot\frac{\langle v_{i},v_{i}\rangle^{2}}{\langle v_{i},1\rangle^{2}}=\frac{\langle v_{i},v_{i}\rangle}{\langle v_{i},1\rangle^{2}}.$
Since the eigenvectors $v_{i}$ are found by solving the equation $D\star_{\mathsf{t}}v_{i}=e_{i}v_{i}$, the components of $v_{i}$ are rational functions of $e_{i}$ with coefficients in the matrix coefficients of $D\star_{\mathsf{t}}$. Therefore, the $\Delta_{i}$ are rational functions in the eigenvalues $e_{i}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$, where the matrix coefficients of $D\star_{\mathsf{t}}$ lie.
To evaluate the term $\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})$, we need only evaluate symmetric rational functions in the derivatives of the eigenvalues $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$. By Proposition 14 of Appendix A, these expressions lie in the field of rational functions of derivatives of the symmetric functions of $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$. The outcome is an explicit calculation of
$\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
The same argument can be used to calculate derivatives
$\nabla_{v_{1}}\nabla_{v_{2}}\cdots\nabla_{v_{k}}\log(\prod_{i}\Delta_{i})\,\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
Next, we consider the term $\nabla_{D^{\star k}}u_{i}\,|_{\mathsf{t}=0}$ of (4.8). Since the eigenvalues of $D^{\star_{\mathsf{t}}k}\star_{\mathsf{t}}$ are simply the $k^{th}$ powers of the eigenvalues of $D\star_{\mathsf{t}}$, we have
$\nabla_{D^{\star k}}u_{i}\,\big{|}_{\mathsf{t}=0}=\Big{(}\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\Big{)}^{k}=e_{i}^{k}\,\big{|}_{\mathsf{t}=0}\,.$
Hence, $\nabla_{D^{\star k}}u_{i}\,|_{\mathsf{t}=0}$ is also determined by the genus 0 theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$.
The difficulty in applying formula (4.8) to calculate $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ lies in controlling the $\mathsf{R}$-matrix terms. We will use our calculation of $\langle D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ to determine $\mathsf{R}^{1}_{ii}\,|_{\mathsf{t}=0}$ up to the nondegeneracy of the Wronskian.
4.4. Proof of Theorem 5
By the divisor equation,
$\langle D,...,D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{{\tiny$\ell$}}}=\left(q\frac{d}{dq}\right)^{\ell-1}\langle D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}\,.$
We define
$\delta_{l}=\langle D,..,D\rangle_{1,\,\framebox{\tiny{$\ell$}}}-\left(q\frac{d}{dq}\right)^{l-1}\frac{1}{48}\nabla_{D}\log\Big{(}\,\prod_{i=1}^{|\mathsf{Part(n)}|}\Delta_{i}\Big{)}\big{|}_{\mathsf{t}=0}.$
By Theorem 1 and the discussion in Section 4.3, $\delta_{l}\in\mathbb{Q}(t_{1},t_{2})(q)$ can be explicitly calculated.
By (4.8), we have, for $\ell\geq 1$,
(4.9) $\begin{split}2\delta_{\ell}&=\left(q\frac{d}{dq}\right)^{\ell-1}\left(\sum_{i=1}^{|\mathsf{Part}(n)|}\mathsf{R}^{1}_{ii}\,\big{|}_{\mathsf{t}=0}\cdot\nabla_{D}u_{i}\,\big{|}_{t=0}\right)\\ &=\sum_{i=1}^{|\mathsf{Part}(n)|}\sum_{k=0}^{\ell-1}\binom{\ell-1}{k}\left(q\frac{d}{dq}\right)^{k}\mathsf{R}^{1}_{ii}\,\big{|}_{\mathsf{t}=0}\cdot\left(q\frac{d}{dq}\right)^{\ell-1-k}\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\,.\end{split}$
By the construction of the $\mathsf{R}$-matrix [14], the derivatives $q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$ are given by
$q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}=\sum_{l}\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}\cdot\left(\nabla_{D}u_{l}\,\big{|}_{\mathsf{t}=0}-\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\right)\cdot\mathsf{R}_{li}^{1}\,\big{|}_{\mathsf{t}=0}\,,$
where the off-diagonal terms $\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}$ are computed by the equation
$\Psi^{-1}\cdot q\frac{d}{dq}\Psi=[\nabla_{D}U,\mathsf{R}^{1}]\,,$
where $\nabla_{D}U$ is the diagonal matrix with diagonal entries $\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}$ and $\Psi$ is the matrix whose columns are normalized eigenvectors. The functions $q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$ are therefore rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2})(q)$, where the matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$ lie. The higher order derivatives
$\left(q\frac{d}{dq}\right)^{k>1}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$
are rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$.
We would like to calculate the diagonal terms $\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$. Define the column vector
$2\vec{\delta}=(2\delta_{1},2\delta_{2},...,2\delta_{|\mathsf{Part}(n)|})^{T}$
of length $|\mathsf{Part}(n)|$ and the column vector
$\vec{r}=\left(\mathsf{R}^{1}_{11}\,\big{|}_{\mathsf{t}=0},...,\mathsf{R}^{1}_{|\mathsf{Part}(n)||\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0},q\frac{d}{dq}\mathsf{R}^{1}_{11}\,\big{|}_{\mathsf{t}=0},...,q\frac{d}{dq}\mathsf{R}^{1}_{|\mathsf{Part}(n)||\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0},\right....,\\[3.0pt] \left.\big{(}q\frac{d}{dq}\big{)}^{|\mathsf{Part}(n)|-1}\mathsf{R}^{1}_{11}\,\big{|}_{\mathsf{t}=0},...,\big{(}q\frac{d}{dq}\big{)}^{|\mathsf{Part}(n)|-1}\mathsf{R}^{1}_{|\mathsf{Part}(n)||\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}\right)^{T}$
of length $|\mathsf{Part}(n)|^{2}$. Equation (4.9) can be written as a matrix equation
(4.10) $2\vec{\delta}=\mathsf{A}\,\vec{r}\,.$
Here, $\mathsf{A}$ is a $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|^{2}$ matrix of the shape
$\mathsf{A}=(\mathsf{W}\,|\,\widehat{\mathsf{W}}),$
where $\mathsf{W}$ is the $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|$ matrix given by
$\mathsf{W}=\left(\begin{array}[]{cccc}\nabla_{D}u_{1}\,\big{|}_{\mathsf{t}=0}&\nabla_{D}u_{2}\,\big{|}_{\mathsf{t}=0}&\ldots&\nabla_{D}u_{|\mathsf{Part}(n)|}\,|_{\mathsf{t}=0}\\[10.0pt] q\frac{d}{dq}\nabla_{D}u_{1}\,\big{|}_{\mathsf{t}=0}&q\frac{d}{dq}\nabla_{D}u_{2}\,\big{|}_{\mathsf{t}=0}&\ldots&q\frac{d}{dq}\nabla_{D}u_{|\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}\\[10.0pt] \vdots&\vdots&\vdots&\vdots\\[10.0pt] (q\frac{d}{dq})^{|\mathsf{Part}(n)|-1}\nabla_{D}u_{1}\,\big{|}_{\mathsf{t}=0}&(q\frac{d}{dq})^{|\mathsf{Part}(n)|-1}\nabla_{D}u_{2}\,\big{|}_{\mathsf{t}=0}&\ldots&(q\frac{d}{dq})^{|\mathsf{Part}(n)|-1}\nabla_{D}u_{|\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}\end{array}\right)$
and $\widehat{\mathsf{W}}$ is an explicit $|\mathsf{Part}(n)|\times(|\mathsf{Part}(n)|^{2}-|\mathsf{Part}(n)|)$ matrix with entries given by
$\binom{\ell-1}{k}\left(q\frac{d}{dq}\right)^{\ell-1-k}\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\,$
for $k>0$. Since $\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}=e_{i}\,\big{|}_{\mathsf{t}=0}$, the matrix coefficients of $\mathsf{A}$ are $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$.
The matrix $\mathsf{W}$ is the Wronskian of the eigenvalues $e_{i}\,|_{\mathsf{t}=0}$ as functions of $\log(q)$. Since the eigenvalues are analytic functions of $\log(q)$, a result of Boecher (see [49, Lemma 1.12]) implies that the eigenvalues are linearly independent over $\mathbb{Q}(t_{1},t_{2})$ if and only if $\det(\mathsf{W})$ is not identically 00.
If $\det(\mathsf{W})$ is nonzero, then $\mathsf{W}$ is invertible. Equation (4.10) then implies
[…]
The $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|^{2}$ matrix $\mathsf{W}^{-1}\mathsf{A}$ is of the form
$\mathsf{W}^{-1}\mathsf{A}=(\mathsf{I}_{|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|}\,|\,\mathsf{W}^{-1}\widehat{\mathsf{W}})\,.$
Via the identity matrix $\mathsf{I}_{|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|}$, equation (4.11) yields equations for
$\mathsf{R}^{1}_{11}\,\big{|}_{\mathsf{t}=0}\,,\,...\,,\,\mathsf{R}^{1}_{|\mathsf{Part}(n)||\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}$
in terms of the functions $\delta_{1},...,\delta_{|\mathsf{Part}(n)|}$, the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,|_{\mathsf{t}=0}$, and the higher $q\frac{d}{dq}$ derivatives of the functions $\mathsf{R}^{1}_{ii}\,|_{\mathsf{t}=0}$.
We have proven that $\mathsf{R}^{1}_{11}\,\big{|}_{\mathsf{t}=0}\,,...,\,\mathsf{R}^{1}_{|\mathsf{Part}(n)||\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}$ can be computed explicitly from
$\Big{\{}\,\langle D,...,D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{{\tiny$\ell$}}}=\left(q\frac{d}{dq}\right)^{\ell-1}\langle D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}\,\Big{\}}_{k=1}^{|\mathsf{Part}(n)|}$
and rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$. ∎
4.5. Proof of Theorem 6
By construction, the final expressions for
$\mathsf{R}^{1}_{11}\,\big{|}_{\mathsf{t}=0}\,,\,...\,,\,\mathsf{R}^{1}_{|\mathsf{Part}(n)||\mathsf{Part}(n)|}\,\big{|}_{\mathsf{t}=0}$
in terms of the $q\frac{d}{dq}$ derivatives of the eigenvalues
$e_{1}\,\big{|}_{\mathsf{t}=0}\,,\,\ldots\,,\,e_{\mathsf{Part}(n)}\,\big{|}_{\mathsf{t}=0}$
with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$ are equivariant under permutations of the indices. Therefore, after substitution in
$\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}=\frac{1}{2}\sum_{i=1}^{|\mathsf{Part}(n)|}\mathsf{R}^{1}_{ii}\,\big{|}_{\mathsf{t}=0}\cdot\nabla_{D^{\star k}}u_{i}\,\big{|}_{\mathsf{t}=0}+\frac{1}{48}\nabla_{D^{\star k}}\log\Big{(}\,\prod_{i=1}^{|\mathsf{Part(n)|}}\Delta_{i}\Big{)}\,\big{|}_{\mathsf{t}=0}\,,$
the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ are symmetric rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$. After an application of Proposition 14 of Appendix A, the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ can be effectively reconstructed from $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ and $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)$. ∎
4.6. The Wronskian
We formulate the following nondegeneracy conjecture for the quantum cohomology of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$.
Conjecture 13.
For all $n\geq 1$, the Wronskian matrix $\mathsf{W}$ associated to $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is nondegenerate:
$\det(\mathsf{W})\neq 0.$
We have verified Conjecture 13 for $n\leq 7$ by computer calculations. As discussed in Section 4.4, Conjecture 13 can reformulated as the assertion that the eigenvalues
$e_{1}\,\big{|}_{\mathsf{t}=0}\,,\,\ldots\,,\,e_{\mathsf{Part}(n)}\,\big{|}_{\mathsf{t}=0}$
are linearly independent over $\mathbb{Q}(t_{1},t_{2})$.
5. Calculations
$\bullet$ For all $n$, the series $\langle(1^{n})\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ has only a constant term in $q$. The calculation was already discussed in Section 1:
(5.1) $\langle(1^{n})\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\text{Coeff}_{Q^{n}}\Big{[}-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\mathcal{P}(Q)\log\mathcal{P}(Q)\Big{]}\,.$
$\bullet$ For all $n$, the series $\langle(2,1^{n-2})\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=-\langle\,D\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is evaluated by Theorem 1,
(5.2) $\langle(2,1^{n-2})\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\left(\mathsf{Tr}_{n}+\sum_{k=2}^{n-1}{\sigma_{-1}(n-k)}\mathsf{Tr}_{k}\right)\,.$
For $2\leq n\leq 5$ and every partition $\mu$ of $n$, we present here closed formulas for the 1-point series
$\langle\,\mu\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
We use a combination of inputs: the full genus 0 theory, the evaluations (5.1) and (5.2), and Getzler’s equation. Once the 1-point series are known, Theorem 4 effectively determines the full genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$.
While Givental’s formula was used in Section 4.4 to prove a structural reconstruction result, calculations are more efficiently obtained from the known series by Getzler’s equation.
$\bullet$ For $n=2$, we have:
$\begin{split}\langle(1,1)\rangle_{1}^{\mathsf{Hilb}^{2}(\mathbb{C}^{2})}&=-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{5}{2}\,,\\[5.0pt] \langle\,(2)\,\rangle_{1}^{\mathsf{Hilb}^{2}(\mathbb{C}^{2})}&=\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{q+1}{q-1}\,.\end{split}$
$\bullet$ For $n=3$, we have:
$\begin{split}\langle(1,1,1)\rangle_{1}^{\mathsf{Hilb}^{3}(\mathbb{C}^{2})}&=-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{29}{6}\,,\\[5.0pt] \langle\,(2,1)\,\rangle_{1}^{\mathsf{Hilb}^{3}(\mathbb{C}^{2})}&=\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{5q^{3}-3q^{2}-3q+5}{(q-1)(q^{2}-q+1)}\,,\\[5.0pt] \langle(3)\rangle_{1}^{\mathsf{Hilb}^{3}(\mathbb{C}^{2})}&=\frac{-1}{12}\frac{(t_{1}+t_{2})}{t_{1}t_{2}}\frac{(t_{1}^{2}+\frac{1}{3}t_{1}t_{2}+t_{2}^{2})(q^{4}+1)-\frac{1}{2}(t_{1}^{2}-\frac{17}{3}t_{1}t_{2}+t_{2}^{2})(q^{3}+q)-(3t_{1}^{2}+13t_{1}t_{2}+3t_{2}^{2})q^{2}}{(q^{2}-q+1)^{2}}\,.\end{split}$
$\bullet$ For $n=4$, we have:
$\begin{split}&\langle(1,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{109}{12}\,,\\[5.0pt] &\langle\,(2,1,1)\,\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{35q^{5}-28q^{4}+23q^{3}+23q^{2}-28q+35}{2(q-1)(q^{2}+1)(q^{2}-q+1)}\,,\end{split}$
$\begin{split}\langle(2,2)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot&\frac{3}{16(q^{2}+1)^{2}(q-1)^{2}}\cdot\\ &\ \left((t_{1}^{2}+\frac{1}{18}t_{1}t_{2}+t_{2}^{2})(q^{6}+1)+(\frac{2}{9}t_{1}^{2}+5t_{1}t_{2}+\frac{2}{9}t_{2}^{2})(q^{5}+q)\right.\\ &\left.+(\frac{11}{9}t_{1}^{2}-\frac{77}{18}t_{1}t_{2}+\frac{11}{9}t_{2}^{2})(q^{4}+q^{2})+(4t_{1}^{2}+\frac{170}{9}t_{1}t_{2}+4t_{2}^{2})q^{3}\right)\,,\end{split}$
$\begin{split}\langle(3,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot&\frac{1}{2(q^{2}-q+1)^{2}(q^{2}+1)^{2}}\cdot\\[5.0pt] &\left((t_{1}^{2}+\frac{23}{36}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{6}t_{1}^{2}-\frac{41}{36}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{7}+q)+(\frac{1}{3}t_{1}^{2}-\frac{257}{36}t_{1}t_{2}+\frac{1}{3}t_{2}^{2})(q^{6}+q^{2})\right.\\ &\left.+(\frac{5}{6}t_{1}^{2}+\frac{137}{12}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{5}+q^{3})-(\frac{8}{3}t_{1}^{2}+\frac{170}{6}t_{1}t_{2}+\frac{8}{3}t_{2}^{2})q^{4}\right)\,,\\[5.0pt] \langle(4)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot&\frac{(q+1)}{4(q^{2}-q+1)(q^{2}+1)^{3}(q-1)}\cdot\\[5.0pt] &\left((t_{1}^{2}-\frac{5}{6}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{3}t_{1}^{2}-\frac{25}{6}t_{1}t_{2}+\frac{5}{3}t_{2}^{2})(q^{7}+q)+(2t_{1}^{2}-13t_{1}t_{2}+2t_{2}^{2})(q^{6}+q^{2})\right.\\[5.0pt] &\left.+(3t_{1}^{2}+\frac{69}{2}t_{1}t_{2}+3t_{2}^{2})(q^{5}+q^{3})-(\frac{10}{3}t_{1}^{2}+39t_{1}t_{2}+\frac{10}{3}t_{2}^{2})q^{4}\right)\,.\end{split}$
$\bullet$ For $n=5$, we have:
$\langle 1\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{907}{60}\,.$
We will write remaining series in terms of the traces [^24] The subscript $m$ of $\mathsf{Tr}_{m}^{\mu}$ is redundant since $m=|\mu|$, but is included for clarity. of quantum multiplication,
$\mathsf{Tr}_{m}^{\mu}=\mathsf{trace}\left(\,\mu\,\star:QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\to QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\,\right)\,.$
The trace which appears in Theorem 1 can be written as
$\mathsf{Tr}_{m}=-\frac{1}{t_{1}+t_{2}}\mathsf{Tr}_{m}^{(2,1^{m-2})}\,.$
Then, we have:
$\begin{split}\langle(2,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=-\frac{t_{1}+t_{2}}{18t_{1}t_{2}}\mathsf{Tr}_{2}^{(2)}-\frac{t_{1}+t_{2}}{16t_{1}t_{2}}\mathsf{Tr}_{3}^{(2,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{4}^{(2,1,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{5}^{(2,1,1,1)}\\ &=\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(\frac{272q^{9}-539q^{8}+760q^{7}-629q^{6}+302q^{5}+302q^{4}-629q^{3}+760q^{2}-539q+272}{6(q-1)(q^{2}+1)(q^{2}-q+1)(q^{4}-q^{3}+q^{2}-q+1)}\right)}}\,,\end{split}$
$\begin{split}\langle(2,2,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{775t_{1}^{2}+733t_{1}t_{2}+775t_{2}^{2}}{1200(t_{1}+t_{2})}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-10t_{1}^{2}-13t_{1}t_{2}-10t_{2}^{2}}{240t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-3}{200(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}+\frac{-25t_{1}^{2}-68t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}\\[5.0pt] &+\frac{50t_{1}^{2}-139t_{1}t_{2}+50t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{-25t_{1}^{2}-44t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,,\end{split}$
$\begin{split}\langle(3,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{175t_{1}^{2}+77t_{1}t_{2}+175t_{2}^{2}}{300(t_{1}+t_{2})}+\frac{-225t_{1}^{2}-532t_{1}t_{2}-225t_{2}^{2}}{3600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{1}{120(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}\\[5.0pt] &+\frac{-3}{100(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{-25t_{1}^{2}-47t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-2}{75(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}\\[5.0pt] &+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,,\end{split}$
$\begin{split}\langle(3,2)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{865t_{1}^{2}+1556t_{1}t_{2}+865t_{2}^{2}}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-158t_{1}^{2}-91t_{1}t_{2}-158t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{20t_{1}^{2}+79t_{1}t_{2}+20t_{2}^{2}}{450(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}+\frac{-1}{300(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}\\[5.0pt] &+\frac{35t_{1}^{2}-77t_{1}t_{2}+35t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-25t_{1}^{2}-77t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{1}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}+\frac{-25t_{1}^{2}-39t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}\\[5.0pt] &+\frac{-23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}+\frac{-4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,,\end{split}$
$\begin{split}\langle(4,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{-430t_{1}^{2}+743t_{1}t_{2}-430t_{2}^{2}}{2400(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-461t_{1}^{2}+503t_{1}t_{2}-461t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{925t_{1}^{2}+1771t_{1}t_{2}+925t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}\\[5.0pt] &+\frac{-100t_{1}^{2}-223t_{1}t_{2}-100t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}+\frac{-113t_{1}t_{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-1}{50(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{-25t_{1}^{2}-63t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}\\[5.0pt] &+\frac{1}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}+\frac{-11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,,\end{split}$
$\begin{split}\langle(5)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{t_{1}t_{2}(110t_{1}^{2}-1391t_{1}t_{2}+110t_{2}^{2})}{600(t_{1}+t_{2})}+\frac{-75t_{1}^{2}-53t_{1}t_{2}-75t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-50t_{1}^{2}+77t_{1}t_{2}-50t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{75t_{1}^{2}+53t_{1}t_{2}+75t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-5t_{1}^{2}-16t_{1}t_{2}-5t_{2}^{2}}{120t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(5)}\\[5.0pt] &+\frac{175t_{1}^{2}+94t_{1}t_{2}+175t_{2}^{2}}{150(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{44t_{1}^{2}+45t_{1}t_{2}+44t_{2}^{2}}{120(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{-60t_{1}^{2}-53t_{1}t_{2}-60t_{2}^{2}}{200(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,1,1)}\\[5.0pt] &+\frac{15t_{1}^{2}+23t_{1}t_{2}+15t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,.\end{split}$
$\bullet$ There is no obstruction (apart from expected nondegeneracies) to extending the above tables to higher $n$. Whether further structures can be found in the $1$-points series $\langle\,\mu\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is an interesting open question.
Appendix A On symmetric functions
Let $\mathbb{K}$ be a field of characteristic 0. Let $\mathbf{z}=(z_{1},\ldots,z_{m})$ be vector of $m$ variables. Let $f_{1}(\mathbf{z}),\ldots,f_{n}(\mathbf{z})$ be $n$ abstract functions, and let
$s_{k}(\mathbf{z})=(-1)^{k}\sum_{I\subseteq[n],|I|=k}\prod_{i\in I}f_{i}$
be the elementary symmetric polynomials in $f_{1},...,f_{n}$. The discriminant is defined by
$\Delta(\mathbf{z})=\prod_{i\neq j}(f_{i}-f_{j})\,.$
We consider the following algebras:
$\bullet$ The standard algebra of symmetric functions,
$\mathsf{Sym}=\mathbb{K}[f_{1},\ldots,f_{n}]^{S_{n}}=\mathbb{K}[s_{1},\ldots,s_{n}]\,,$
is defined by taking invariants of the $S_{n}$-action on $\mathbb{K}[f_{1},\ldots,f_{n}]$. The $S_{n}$-action is defined by permuting the indices of $f_{i}$. By construction, $\Delta\in\mathsf{Sym}$.
$\bullet$ Let $\mathsf{Df}=\big{\{}\frac{\partial^{a_{1}+\ldots+a_{m}}}{\partial z_{1}^{a_{1}}\cdots\partial z_{m}^{a_{m}}}f_{i}\big{\}}_{(a_{1},\ldots,a_{m})\in(\mathbb{Z}_{\geq 0})^{m}}$ be the set of all partial derivatives of the functions $f_{1},\ldots,f_{n}$,
$\mathsf{Df}=\big{\{}f_{1},\ldots,f_{n},\ldots,\frac{\partial f_{i}}{\partial z_{j}},\ldots,\frac{\partial^{2}f_{i}}{\partial z_{j_{1}}\partial z_{j_{2}}},\ldots\big{\}}\,.$
[…]
$\mathsf{SymDf}=\mathbb{K}[\mathsf{Df}]^{S_{n}}\,.$
$\bullet$ Let $\mathsf{Ds}=\big{\{}\frac{\partial^{a_{1}+\ldots+a_{m}}}{\partial z_{1}^{a_{1}}\cdots\partial z_{m}^{a_{m}}}s_{i}\big{\}}_{(a_{1},\ldots,a_{m})\in(\mathbb{Z}_{\geq 0})^{m}}$ be the set of all partial derivatives of the elementary symmetric functions $s_{1},\ldots,s_{n}$. Let $\mathbb{K}[\mathsf{Ds}]$ be the algebra of polynomials in the functions $\mathsf{Ds}$.
We present a proof of the following result (which is likely known to experts, but we were unable to find a reference).
Proposition 14.
$\mathsf{SymDf}\subseteq\mathbb{K}[\mathsf{Ds}][1/\Delta]$
Proof.
Let ${\mathbf{a}}=(a_{1},\ldots,b_{m})$ and ${\mathbf{b}}=(b_{1},\ldots,b_{m})$ both be elements of $(\mathbb{Z}_{\geq 0})^{m}$. We define
  1. (i)
    $\mathbf{a}\leq\mathbf{b}$ if $a_{j}\leq b_{j}$ for all $1\leq j\leq m$,
  2. (ii)
    $\mathbf{a}<\mathbf{b}$ if $\mathbf{a}\leq\mathbf{b}$ and $a_{j}<b_{j}$ for some $j$.
Let $\mathsf{Ds}_{\leq\mathsf{b}}=\big{\{}\frac{\partial^{a_{1}+\ldots+a_{m}}}{\partial z_{1}^{a_{1}}\cdots\partial z_{m}^{a_{m}}}s_{i}\big{\}}_{\mathsf{a}\leq\mathsf{b}}$ be a finite set of partial derivatives of the elementary symmetric functions $s_{1},\ldots,s_{n}$. Similarly, let $\mathsf{DY}_{<\mathsf{b}}=\big{\{}\frac{\partial^{a_{1}+\ldots+a_{m}}}{\partial z_{1}^{a_{1}}\cdots\partial z_{m}^{a_{m}}}Y\big{\}}_{\mathsf{a}<\mathsf{b}}$ be the finite set of partial derivatives of a single abstract function $Y(\mathbf{z})$.
Consider the polynomial
$P(x)=x^{n}+s_{1}x^{n-1}+\ldots+s_{n}=\prod_{i}(x-f_{i})\,.$
We can take the $\frac{\partial}{\partial z_{j}}$ derivative of the relation $P(f_{i})=0$ :
(A.1) $\frac{\partial f_{i}}{\partial z_{j}}P_{x}(f_{i})+\frac{\partial s_{1}}{\partial z_{j}}f_{i}^{n-1}+\ldots+\frac{\partial s_{n-1}}{\partial z_{j}}f_{i}+\frac{\partial s_{n}}{\partial z_{j}}=0\,,$
where $P_{x}$ is the derivative of $P$ as a polynomial in $x$. By a simple calculation,
$\Delta=\prod_{i=1}^{n}P_{x}(f_{i})\,.$
Let $\mathbf{b}\in(\mathbb{Z}_{\geq 0})^{m}$. By repeatedly taking derivatives of (A.1), we find that there is a universal polynomial
$\Phi_{\mathbf{b}}(\mathsf{Ds}_{\leq\mathbf{b}},\mathsf{DY}_{<\mathbf{b}})\in\mathbb{K}[\mathsf{Ds}_{\leq\mathbf{b}},\mathsf{DY}_{<\mathbf{b}}]$
which satisfies the following property: for every $f_{i}$,
(A.2) $\frac{\partial^{b_{1}+\ldots+b_{m}}f_{i}}{\partial z_{1}^{b_{1}}\cdots\partial z_{m}^{b_{m}}}P_{x}(f_{i})=\Phi_{\mathbf{b}}|_{Y=f_{i}}\,.$
Therefore, there is a universal polynomial $\Omega_{\mathbf{b}}(\mathsf{Ds}_{\leq\mathbf{b}},Y)\in\mathbb{K}[\mathsf{Ds}_{\leq\mathbf{b}},Y]$ for which
$\frac{\partial^{b_{1}+\ldots+b_{m}}f_{i}}{\partial z_{1}^{b_{1}}\cdots\partial z_{m}^{b_{m}}}P_{x}(f_{i})^{N_{\mathbf{b}}}=\Omega_{\mathbf{b}}(\mathsf{Ds}_{\leq\mathbf{b}},f_{i})\,$
for a (possibly large) integer $N_{\mathbf{b}}$.
We now take an arbitrary monomial in the functions of $\mathsf{Df}$:
$M=\prod_{i=1}^{n}\prod_{u=1}^{v_{i}}\partial_{\mathbf{b}(i,u)}f_{i}\,,$
where we have used the notation
$\partial_{\mathbf{b}(i,u)}f_{i}=\frac{\partial^{b_{1}(i,u)+\ldots+b_{m}(i,u)}f_{i}}{\partial z_{1}^{b_{1}(i,u)}\cdots\partial z_{m}^{b_{m}(i,u)}}\,.$
By (A.2), we have
(A.3) $\Delta^{\sum_{i=1}^{n}\sum_{u=1}^{v_{i}}N_{\mathbf{b}(i,u)}}\cdot M=\prod_{i=1}^{n}\prod_{u=1}^{v_{i}}\Omega_{\mathbf{b}(i,u)}(\mathsf{Ds}_{\leq\mathbf{b}(i,u)},f_{i})\cdot\Delta_{i}^{N_{\mathbf{b}(i,u)}}\,,$
where $\Delta_{i}=\Delta/P_{x}(f_{i})$.
Consider next the $S_{n}$-invariant element
$\mathsf{sym}M=\sum_{\sigma\in S_{n}}\sigma(M)\in\mathsf{SymDf}\,.$
Using (A.3), we obtain
(A.4) $\mathsf{sym}M=\Delta^{-\sum_{i=1}^{n}\sum_{u=1}^{v_{i}}N_{\mathbf{b}(i,u)}}\sum_{\sigma\in S_{n}}\sigma\left(\prod_{i=1}^{n}\prod_{u=1}^{v_{i}}\Omega_{\mathbf{b}(i,u)}(\mathsf{Ds}_{\leq\mathbf{b}(i,u)},f_{i})\cdot\Delta_{i}^{N_{\mathbf{b}(i,u)}}\right)\,$
where the right side of (A.4) lies in $\mathbb{K}[\mathsf{Ds}][1/\Delta]$. As $M$ varies over all monomials in $\mathbb{K}[\mathsf{Df}]$, the elements $\mathsf{sym}M$ generate $\mathsf{SymDf}$. ∎

### v2

@@
 We calculate the genus 1 Gromov-Witten theory of the Hilbert scheme $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ of points in the plane.
-The fundamental 1-point invariant (with a divisor insertion) is calculated using a correspondence with the families local curve Gromov-Witten theory over the moduli space ¯$\overline{\mathcal{M}}_{1,1}$.
+The fundamental 1-point invariant (with a divisor insertion) is calculated using a correspondence with the families local curve Gromov-Witten theory over the moduli space $\overline{\mathcal{M}}_{1,1}$.
 The answer exactly matches a parallel calculation related to the Noether-Lefschetz geometry of the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties.
@@
 $\dim_{\mathbb{C}}{\mathbb{C}[x,y]}/{\mathcal{I}}=n\,.$
-An open dense set of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ parameterizes ideals associated to configurations of $n$ distinct unordered points.
+An open dense subset of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ parameterizes ideals associated to configurations of $n$ distinct unordered points.
 The geometry of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ has been studied from many points of view for several decades now, see [12, 17, 18, 21, 34].
@@
 $\frac{1}{\Pi_{i}\mu_{i}}[V_{\mu}]$
-where $[V_{\mu}]$ is (the cohomological dual of) the class of the subvariety of $\mathsf{Hilb}^{|\mu|}(\mathbb{C}^{2})$ with generic element given by a union of schemes of lengths
+where $V_{\mu}]$ is (the cohomological dual of) the class of the subvariety of $\mathsf{Hilb}^{|\mu|}(\mathbb{C}^{2})$ with generic element given by a union of schemes of lengths
 $\mu_{1},\ldots,\mu_{\ell(\mu)}$
@@
 (0.1) $\mathcal{O}/\mathcal{I}\rightarrow\mathsf{Hilb}^{n}(\mathbb{C}^{2})\,$,
-with fiber $\mathbb{C}[x,y]/\mathcal{I}$ over $[\mathcal{I}]\in\mathsf{Hilb}^{n}(\mathbb{C}^{2})$.
+with fiber $\mathbb{C}[x,y]/\mathcal{I}$ over $\mathcal{I}]\in\mathsf{Hilb}^{n}(\mathbb{C}^{2})$.
 The $\mathsf{T}$-action on $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ lifts canonically to the tautological bundle (0.1).
@@
 see [26].
+0.2. Quantum cohomology
 The $\mathsf{T}$-equivariant quantum cohomology of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ has been determined in [39].
@@
 The matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$ lie in the field of rational functions in $q$ (with coefficients [^6] In the context of quantum cohomology, the definitions require localization, so we will always consider $H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ and $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as modules over $\mathbb{Q}(t_{1},t_{2})$.
-in $\mathbb{Q}(t_{1},t_{2})$) and determine all genus 00 Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ by [37, Section 4.2].
+in $\mathbb{Q}(t_{1},t_{2})$) and determine all genus 0 Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ by [39, Section 4.2].
 We are interested here in the $\mathsf{T}$-equivariant Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ in genus 1. Let $\mu^{1},\ldots,\mu^{r}$ be partitions of $n$. Define
 (0.2) $\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\,=\,\sum_{d=0}^{\infty}\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1,d}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}q^{d}\,\in\mathbb{Q}(t_{1},t_{2})[[q]].$
-The series (0.2) is always a rational function in $q$,
-$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2})(q).\,$
+The series (0.2) is always a rational [^7] There is a gap in the proof of rationality in the published paper [46] which is filled in a revision available on the authors’ websites.
+The gap does not affect the proof in [46] of the triangle of correspondences [46, Theorem 2] in the form of analytic continuation from $q=0$ to q=1q=-1, which is all that is needed here.
+function [46] in $q$,
+$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2},q)\,.\,$
 The first nontrivial computation in genus 1 appeared in [46]:
@@
 Our goal here is to provide calculations of all of the series (0.2) starting with the basic case of $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$.
+The genus 0 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is a very rich subject with connections to the quantum Calogero-Sutherland many body system [39], to the crepant resolution conjecture [3], and to quantum groups [30].
+The question of computing $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ for all $n$ was posed in [44, 46].
+We will see that the genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ provides new connections: the theory is fundamentally linked to the geometry of the moduli space of abelian varieties.
+There are several natural open questions.
+What is the structure of the genus 1 Gromov-Witten theory of more general Nakajima quiver varieties?
+Is there an exact calculation of the full genus Gromov-Witten theory?
 0.3. Gromov-Witten theory of families of local elliptic curves
-Let ¯$\overline{\mathcal{M}}_{1,r}$ be the moduli space of Deligne-Mumford stable curves of genus $1$ with $r$ markings. [^7] We will always assume $r>0$ for stability. Let
+Let $\overline{\mathcal{M}}_{1,r}$ be the moduli space of Deligne-Mumford stable curves of genus $1$ with $r$ markings. [^8] We will always assume $r>0$ for stability. Let
 $\pi:\mathcal{E}\rightarrow\overline{\mathcal{M}}_{1,r}\,$
@@
 $\pi_{\mathbb{C}^{2}}:\mathcal{E}\times\mathbb{C}^{2}\rightarrow\overline{\mathcal{M}}_{1,r}$
-be the universal local curve over ¯$\overline{\mathcal{M}}_{1,r}$. The torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ acts on the $\mathbb{C}^{2}$ factor as before.
-Let $\mu^{1},\ldots,\mu^{r}\in\mathrm{Part}(n)$, and let ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ be the moduli space of stable [^8] The superscript $\bullet$ indicates possibly disconnected domain curves (but no connected component of the domain is contracted to a point).
+be the universal local curve over $\overline{\mathcal{M}}_{1,r}$. The torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ acts on the $\mathbb{C}^{2}$ factor as before.
+Let $\mu^{1},\ldots,\mu^{r}\in\mathrm{Part}(n)$, and let $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ be the moduli space of stable [^9] The superscript $\bullet$ indicates possibly disconnected domain curves (but no connected component of the domain is contracted to a point).
 We follow the conventions of [4].
@@
 condition $\mu^{i}$ over the divisor $p_{i}\times\mathbb{C}^{2}$.
-Since the degree $n$ is recorded in the size of the partitions $\mu^{i}$, we omit $n$ from the notation for ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$.
-The moduli space ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ has $\pi_{\mathbb{C}^{2}}$-relative virtual dimension
-$-nr+\sum_{i=1}^{r}\ell(\mu^{i})\,.$
+Since the degree $n$ is recorded in the size of the partitions $\mu^{i}$, we omit $n$ from the notation for $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$.
+The moduli space $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ has $\pi_{\mathbb{C}^{2}}$-relative virtual dimension
+$nr+\sum_{i=1}^{r}\ell(\mu^{i})\,.$
 The Gromov-Witten series of the family $\pi_{\mathbb{C}^{2}}$ is defined by
@@
 $2g[b]-2=b+nr-\sum_{i=1}^{r}\ell(\mu^{i})\,.$
-The moduli space of stable maps ¯$\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}$ is empty unless $g[b]$ is an integer.
-The virtual class [¯$\Big{[}\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}$ is the $\pi_{\mathbb{C}^{2}}$-relative $\mathsf{T}$-equivariant virtual class of the family of relative stable maps to the fibers of $\pi_{\mathbb{C}^{2}}$.
+The moduli space of stable maps $\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}$ is empty unless $g[b]$ is an integer.
+The virtual class $\Big{[}\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}$ is the $\pi_{\mathbb{C}^{2}}$-relative $\mathsf{T}$-equivariant virtual class of the family of relative stable maps to the fibers of $\pi_{\mathbb{C}^{2}}$.
 We define
 $\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{g}^{\pi_{\mathbb{C}^{2}},\bullet}$ $=$ $\text{Coeff}_{u^{2g-2}}\Big{[}\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}^{\pi_{\mathbb{C}^{2}},\bullet}\Big{]}$
-$=$ ¯$\int_{\overline{\mathcal{M}}_{1,r}}\epsilon_{*}\left(\Big{[}\overline{\mathcal{M}}^{\bullet}_{g}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}\right).$
+$=$ $\int_{\overline{\mathcal{M}}_{1,r}}\epsilon_{*}\left(\Big{[}\overline{\mathcal{M}}^{\bullet}_{g}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}\right).$
 To emphasize the degree $n$, we will sometimes use the notation
@@
 $\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}=(-i)^{\sum_{i=1}^{r}\ell(\mu^{i})-|\mu^{i}|}\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}^{\pi_{\mathbb{C}^{2}},\bullet}$
-after the variable change $-q=e^{iu}$.
+after the variable change $q=e^{iu}$.
 The calculation of the Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ in genus 1 is therefore equivalent to the calculation of the Gromov-Witten invariants of the family $\pi_{\mathbb{C}^{2}}$ of local elliptic curves.
@@
 $\nu:\mathcal{X}_{g}\rightarrow\mathcal{A}_{g}$
-be the universal principally polarized abelian variety. We refer the reader to [2] for the foundations of the study of the moduli of abelian varieties.
+be the universal family of principally polarized abelian varieties. We refer the reader to [2] for the foundations of the study of the moduli of abelian varieties.
 A general abelian variety $(X,\theta)$ parameterized by $\mathcal{A}_{g}$ has Picard number $1$.
 The Noether-Lefschetz locus of $\mathcal{A}_{g}$ parameterizes abelian varieties with Picard number at least 2.
-The simplest components of the Noether-Lefschetz locus of $\mathcal{A}_{g}$ are related to the geometry of elliptic curves on abelian varieties.
+The simplest irreducible components of the Noether-Lefschetz locus of $\mathcal{A}_{g}$ are related to the geometry of elliptic curves on abelian varieties.
 For $g\geq 1$ and $n\geq 1$, let
-$\mathsf{NL}_{g,n}=\left\{(X,\theta)\in\mathcal{A}_{g}\left|\begin{array}[]{c}X\text{ contains a subgroup }E\subset X\\ \text{ which is an elliptic curve of degree }\theta\cdot[E]=n\end{array}\right.\right\}.$
-Let $[{\mathsf{NL}}_{g,n}]\in\mathsf{CH}^{g-1}(\mathcal{A}_{g})$ be the associated cycle class.101010 The $g=1$ case is degenerate: $[\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$ by definition.
+[…]
+Let ${\mathsf{NL}}_{g,n}]\in\mathsf{CH}^{g-1}(\mathcal{A}_{g})$ be the associated cycle class.111111 The $g=1$ case is degenerate: $\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$ by definition.
 All Chow groups are taken with $\mathbb{Q}$-coefficients.
 The following linear combination of components plays a geometrically important role [^12] As usual, $\sigma_{r}(n)=\sum_{d|n}d^{r}$.
-$[\widetilde{\mathsf{NL}}_{g,n}]=\sum_{n^{\prime}|n}\sigma_{1}\left(\frac{n}{n^{\prime}}\right)[\mathsf{NL}_{g,n^{\prime}}]\ \in\ \mathsf{CH}^{g-1}(\mathcal{A}_{g})\,,$
+$\widetilde{\mathsf{NL}}_{g,n}]=\sum_{n^{\prime}|n}\sigma_{1}\left(\frac{n}{n^{\prime}}\right)[\mathsf{NL}_{g,n^{\prime}}]\ \in\ \mathsf{CH}^{g-1}(\mathcal{A}_{g})\,$,
 see [20].
@@
 $\mathsf{taut}([\mathsf{NL}_{g,n}])=\frac{n^{2g-1}g}{6|B_{2g}|}\prod_{p|n}(1-p^{2-2g})\lambda_{g-1}$
-or, equivalently [^12] Equation (0.4) is correct also for $g=1$ with the convention $[\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$.,
+or, equivalently [^13] Equation (0.4) is correct also for $g=1$ with the convention $\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$.,
 (0.4) $\mathsf{taut}\left(\frac{(-1)^{g}}{24}\lambda_{g-1}+\sum_{n=1}^{\infty}[\widetilde{\mathsf{NL}}_{g,n}]\,Q^{n}\right)=\frac{(-1)^{g}}{24}E_{2g}(Q)\lambda_{g-1},$
@@
 $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\frac{5q^{3}-3q^{2}-3q+5}{(q-1)(q^{2}-q+1)}\right)$
-$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(-5-7q-q^{2}+2q^{3}-q^{4}-7q^{5}-10q^{6}-7q^{7}+...\right),$
+[…]
 $\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}$ $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\mathsf{Tr}_{4}+\mathsf{Tr}_{3}+\frac{3}{2}\mathsf{Tr}_{2}\right)$
 $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(\frac{35q^{5}-28q^{4}+23q^{3}+23q^{2}-28q+35}{2(q-1)(q^{2}+1)(q^{2}-q+1)}\right)}}$
-$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(-\frac{35}{2}-21q-q^{2}-3q^{3}-17q^{4}-21q^{5}-19q^{6}-21q^{7}+...\right)\,,$
+[…]
 $\big{\langle}\,D\,\big{\rangle}_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}$ $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\mathsf{Tr}_{5}+\mathsf{Tr}_{4}+\frac{3}{2}\mathsf{Tr}_{3}+\frac{4}{3}\mathsf{Tr}_{2}\right)$
 $=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(-\frac{272q^{9}-539q^{8}+760q^{7}-629q^{6}+302q^{5}+302q^{4}-629q^{3}+760q^{2}-539q+272}{6(q-1)(q^{2}+1)(q^{2}-q+1)(q^{4}-q^{3}+q^{2}-q+1)}\right)}}$
-$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\frac{136}{3}+\frac{277}{6}q-\frac{41}{6}q^{2}+\frac{17}{3}q^{3}+\frac{151}{6}q^{4}+\frac{127}{6}q^{5}+\frac{101}{3}q^{6}+\frac{277}{6}q^{7}+...\right).$
+[…]
 0.5.2. Families of local elliptic curves
@@
 for stable maps with connected domains.
-Let ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ be the moduli space of stable [^13] The superscript $\circ$ indicates connected domain curves. relative maps to the fibers of $\pi$,
+Let $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ be the moduli space of stable [^14] The superscript $\circ$ indicates connected domain curves. relative maps to the fibers of $\pi$,
 $\epsilon:\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)\rightarrow\overline{\mathcal{M}}_{1,1}\,.$
@@
 is the moduli space of stable maps of 1-pointed connected genus $g$ curves to $E$ of degree $n$ relative to the divisors determined by the nodes of $C$.
-The moduli space ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ has $\pi$-relative virtual dimension $12g-$ 1 and total virtual dimension $2g$.
+The moduli space $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ has $\pi$-relative virtual dimension 2g12g-1 and total virtual dimension $2g$.
 For $g\geq 2$, let
@@
 $\mathsf{taut}:\mathsf{CH}^{*}(\mathcal{A}_{g})\rightarrow\mathsf{R}^{*}(\mathcal{A}_{g})\,.$
-Let $\mathcal{M}_{g}^{\mathsf{ct}}$ be the $33g-$ 3 dimensional moduli space of genus $g$ curves of compact type, and let
+Let $\mathcal{M}_{g}^{\mathsf{ct}}$ be the 3g33g-3 dimensional moduli space of genus $g$ curves of compact type, and let
 $\mathsf{Tor}:\mathcal{M}_{g}^{\mathsf{ct}}\rightarrow\mathcal{A}_{g}$
@@
 0.6. Reduction of multi-point series to 1-point series
-By [37, Section 4.2], the quantum powers of $D$ generate the quantum cohomology $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$. More precisely, the set
+By [39, Section 4.2], the quantum powers of $D$ generate the quantum cohomology $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$. More precisely, the set
 (0.7) $\{1,D,D^{*2},D^{*3},\ldots,D^{*(|\mathsf{Part}(n)|-1)}\}$
-is a basis of $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as a $\mathbb{Q}(t_{1},t_{2})(q)$-vector space.
+is a basis of $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as a $\mathbb{Q}(t_{1},t_{2},q)$-vector space.
 Our first reduction result is that the genus 1 Gromov-Witten theory (0.2) of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ can be effectively reduced to 1-point series in the basis (0.7).
@@
 To every genus $1$ series $\big{\langle}D^{*k_{1}},...,D^{*k_{\ell}}\big{\rangle}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$, there are canonically associated functions
-[…]
+(0.8) $\big{\{}C_{k,m}\big{\}}_{0\leq k\leq|\mathsf{Part}(n)|-1\,,\,0\leq m\leq\ell-1}\subset\mathbb{Q}(t_{1},t_{2},q)$
 for which the following equation holds:
@@
 by constant map invariants in [46].
-The resulting control of the full Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ was used in [44] to prove both the crepant resolution and the GW/DT correspondences associated to the geometry, but was not sufficient to obtain closed form evaluations of any higher genus series (because of the difficulty in the determination and the lack of closed forms for the $\mathsf{R}$-matrix).
+The resulting control of the full Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ was used in [46] to prove both the crepant resolution and the GW/DT correspondences associated to the geometry, but was not sufficient to obtain closed form evaluations of any higher genus series (because of the lack of closed forms for the $\mathsf{R}$-matrix).
 In genus 1, the Givental-Teleman reconstruction formula reduces to an earlier equation due to Givental [15] on the associated Frobenius manifold:
@@
 1. Degree 0 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$
-We expand here explicitly the formula of Theorem 1 for the degree 00 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to show the connections to Hilbert scheme calculations of Carlsson and Okounkov [6].
+We expand here explicitly the formula of Theorem 1 for the degree 0 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to show the connections to Hilbert scheme calculations of Carlsson and Okounkov [7].
 Consider first the generating series over the Hilbert schemes of points:
@@
 The geometric formula for the $q=0$ term of $\langle D\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is obtained via identification of the virtual fundamental class:
-[¯$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(D)$ $=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
+$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(D)$ $=$ $\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
 $=$ $-\frac{1}{24}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})\,.$
@@
 (1.3) $\sum_{n=0}^{\infty}Q^{n}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})=\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}(1+\log\mathcal{P}(Q))\cdot\frac{1}{2}\mathcal{P}(Q)(\mathcal{E}_{2}(Q)-\mathcal{E}_{3}(Q))\,.$
-Consider now the series $\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}$ of [6, Corollary 3] where we take the line bundle $\mathcal{L}$ to be $\mathcal{O}$ with equivariant weight $m$ as in [6, Section 2.1.1]:
+Consider now the series $\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}$ of [7, Corollary 3] where we take the line bundle $\mathcal{L}$ to be $\mathcal{O}$ with equivariant weight $m$ as in [7, Section 2.1.1]:
 $\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}$ $=$ $\sum_{n=0}^{\infty}Q^{n}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})},m)$
 $=$ $\sum_{n=0}^{\infty}Q^{n}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot\left(c_{\mathsf{top}}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})+m\cdot c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})+m^{2}\cdot c_{\mathsf{top}-2}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})\ldots\right)\,.$
-The left side of (1.3) is the coefficient of $m^{1}$ of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$. By [6, Corollary 3] together with evaluations from [6, Section 2.2.2], we have
+The left side of (1.3) is the coefficient of $m^{1}$ of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$. By [7, Corollary 3] together with evaluations from [7, Section 2.2.2], we have
 $\frac{\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}}{\big{\langle}1\big{\rangle}}=\frac{1}{2}(\mathcal{E}_{2}(Q)-\mathcal{E}_{3}(Q))\cdot\frac{(t_{1}+t_{2})(t_{1}+m)(t_{2}+m)}{t_{1}t_{2}}\,$,
@@
 $=$ $\prod_{n\geq 1}(1-Q^{n})^{(\frac{m(-t_{1}-t_{2}-m)}{t_{1}t_{2}}-1)}\,$
-by [6, Corollary 1]. The matching (1.3) then follows from a simple algebraic expansion of the $m^{1}$ coefficient of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$.
+by [7, Corollary 1]. The matching (1.3) then follows from a simple algebraic expansion of the $m^{1}$ coefficient of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$.
 The degree 0 term of Gromov-Witten series $\big{\langle}1\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ can be studied similarly.
 The geometric formula for the $q=0$ term of $\langle 1\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ is
-[¯$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(1)$ $=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
+$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(1)$ $=$ $\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
 $=$ $-\frac{1}{24}\int_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c_{\mathsf{top}-1}(T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})\,$,
-which equals the $m^{1}$ coefficient of $-\frac{1}{24}\big{\langle}1\big{\rangle}$. A calculation then yields
+which equals the $m^{1}$ coefficient of $\frac{1}{24}\big{\langle}1\big{\rangle}$. A calculation then yields
 $\text{Coeff}_{m^{1}}\Big{[}\langle 1\rangle\Big{]}=\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\mathcal{P}(Q)\log\mathcal{P}(Q)\,.$
@@
 2.2. The $n=0$ case
-The $n=0$ invariant concerns the moduli space of degree 0 stable maps ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)$. After imposing the evaluation condition on the marking, the moduli space is
-¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)\supset\text{ev}_{1}^{-1}(\mathsf{p}_{1})=\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}\,$
-with obstruction bundle $\mathbb{E}_{g}^{\vee}\otimes\mathsf{T}$, where $\mathsf{T}$ is the tangent line on ¯$\overline{\mathcal{M}}_{1,1}$ associated to the marking.
+The $n=0$ invariant concerns the moduli space of degree 0 stable maps $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)$. After imposing the evaluation condition on the marking, the moduli space is
+$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)\supset\text{ev}_{1}^{-1}(\mathsf{p}_{1})=\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}\,$
+with obstruction bundle $\mathbb{E}_{g}^{\vee}\otimes\mathbb{T}$, where $\mathbb{T}$ is the tangent line on $\overline{\mathcal{M}}_{1,1}$ associated to the marking.
 $\bullet$ For $g\geq 2$, we evaluate the $n=0$ terms by:
-$\big{\langle}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\big{\rangle}_{g,0}^{\pi,\circ}$ $=$ [¯$\int_{[\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}\lambda_{g}\lambda_{g-2}$
-$=$ ¯$\int_{\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\lambda_{g}\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{T})$
+$\big{\langle}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\big{\rangle}_{g,0}^{\pi,\circ}$ $=$ $\int_{[\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}\lambda_{g}\lambda_{g-2}$
+$=$ $\int_{\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\lambda_{g}\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathbb{T})$
 $=$ $(-1)^{g}\int_{\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\lambda_{g}\lambda_{g-2}(\lambda_{g}+\lambda_{g-1}\boxtimes\psi)$
@@
 $=$ $\frac{(-1)^{g}}{24}\frac{|B_{2g}|}{4g}\frac{|B_{2g-2}|}{(2g-2)!}\,.$
-Here, $\psi_{1}$ denotes the cotangent line at the marking of ¯$\overline{\mathcal{M}}_{g,1}$.
-In the third equality, we have denoted the dual of $c_{1}(\mathsf{T})$ by $\psi$, the cotangent line at the marking of ¯$\overline{\mathcal{M}}_{1,1}$.
-In the fourth equality, we have used the dilaton equation and the Hodge integral evaluated in [9, Theorem 4].
+Here, $\psi_{1}$ denotes the cotangent line at the marking of $\overline{\mathcal{M}}_{g,1}$.
+In the third equality, we have denoted the dual of $c_{1}(\mathbb{T})$ by $\psi$, the cotangent line at the marking of $\overline{\mathcal{M}}_{1,1}$.
+In the fourth equality, we have used the dilaton equation and the Hodge integral evaluated in [10, Theorem 4].
 The final evaluation agrees with Theorem 2 since the Eisenstein series start with 1,
@@
 $\bullet$ For $g=1$, we evaluate the $n=0$ term by:
-$\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,0}^{\pi,\circ}$ $=$ [¯$\int_{[\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}$
-$=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\cdot\mathsf{e}(\mathbb{E}_{1}^{\vee}\otimes\mathsf{T})$
+$\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,0}^{\pi,\circ}$ $=$ $\int_{[\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}$
+$=$ $\int_{\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\cdot\mathsf{e}(\mathbb{E}_{1}^{\vee}\otimes\mathbb{T})$
 $=$ $(-1)\int_{\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}(\lambda_{1}+\lambda_{0}\boxtimes\psi)$
@@
 For $g=1$ and $n\geq 1$, we can evaluate the integral $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ by hand.
-There are no branch points, so the cotangent line on the domain is pulled-back from the cotangent line of ¯$\overline{\mathcal{M}}_{1,1}$.
+There are no branch points, so the cotangent line on the domain is pulled-back from the cotangent line of $\overline{\mathcal{M}}_{1,1}$.
 Using the latter cotangent line class, we can express $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ as $\frac{1}{24}$ times a Gromov-Witten invariant of maps to a fixed elliptic target $(E,p_{1})$:
@@
 induced by $\pi_{S}$ is of degree $48$. Therefore,
-[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{48}\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}.$
+$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{48}\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}.$
 The moduli space of stable maps to the fibers of $\pi_{S}:S\rightarrow\mathbb{P}^{1}$ lies over $\mathbb{P}^{1}$,
@@
 The following vanishing holds for $g\geq 2$ and $n>0$:
-[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)=0\,.$
+$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)=0\,.$
 The motivation for Proposition 7 comes from the vanishing of Gromov-Witten invariants for $K3$ surfaces in non-zero curve classes.
@@
 Proposition 7 allow us to exchange the families Hodge integral
-[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}$
+$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}$
 for a fixed target Gromov-Witten invariant. We first expand the Euler class as
 (2.2) $\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)=(-1)^{g}\lambda_{g}+(-1)^{g-1}\lambda_{g-1}\cdot\epsilon_{S}^{*}(2[\mathsf{pt}])\,$,
-where $[\mathsf{pt}]\in\mathsf{CH}_{0}(\mathbb{P}^{1})$ is the point class. By Proposition 7, we obtain
-[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\Big{(}(-1)^{g}\lambda_{g}+(-1)^{g-1}\lambda_{g-1}\cdot 2\epsilon_{S}^{*}([\mathsf{pt}]))\Big{)}=0\,.$
+where $\mathsf{pt}]\in\mathsf{CH}_{0}(\mathbb{P}^{1})$ is the point class. By Proposition 7, we obtain
+$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\Big{(}(-1)^{g}\lambda_{g}+(-1)^{g-1}\lambda_{g-1}\cdot 2\epsilon_{S}^{*}([\mathsf{pt}]))\Big{)}=0\,.$
 Hence,
-[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g}\lambda_{g-2}$ $=$ $2\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-1}\lambda_{g-2}\cdot[\mathsf{pt}]$
+$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g}\lambda_{g-2}$ $=$ $2\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-1}\lambda_{g-2}\cdot[\mathsf{pt}]$
 $=$ $2\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(E,n)]^{vir}}\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}\,$,
@@
 The evaluation of the required integral for $(E,p_{1})$ follows from the methods of [36],
-(2.3) [¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(E,n)]^{vir}}\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}=\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,,$
+(2.3) $\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(E,n)]^{vir}}\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}=\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,$,
 as we will explain in Section 2.6. The integral can also be obtained using the study of the Gromov-Witten theory of target curves [37] and was first calculated by Pixton in [48].
 Theorem 2 then follows from
-[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{24}\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,$
+$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{24}\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,$
 and the definition of the Eisenstein series (2.1). ∎
@@
 The universal curve $\mu:\mathcal{C}\rightarrow\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ carries a universal evaluation map
-$\mathcal{C}$$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$$f$$\mu$$\nu$
+[…]
 to the universally expanded target
-$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\times_{\mathbb{P}^{1}}S$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$h$$\nu$$\pi_{M}$
-The target $\mathcal{T}$ is a family of elliptic curves over ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ with possible expansion over the 24 nodal fibers $R_{i}$.
+[…]
+The target $\mathcal{T}$ is a family of elliptic curves over $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ with possible expansion over the 24 nodal fibers $R_{i}$.
 The only permitted expansions over $R_{i}$ are simple circuits of rational curves.
@@
 Consider next the stabilization map
-$\mathcal{C}$$\mathcal{C}_{\mathsf{st}}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$\mathsf{st}$$\mu$$\mu_{\mathsf{st}}$
+$\mathcal{C}$$\mathcal{C}_{\mathsf{st}}$$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$\mathsf{st}$$\mu$$\mu_{\mathsf{st}}$
 By the geometry of relative stable maps, the contraction $\mathsf{st}$ is an isomorphism (except for collapsing chains of unstable rational curves to nodes). Therefore,
@@
 and we obtain $\mu_{*}\omega_{\mu}\cong\mathbb{E}_{g}$.
-Since the moduli space ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ parametrizes relative maps to the fibers of $\pi_{S}$, we have a pull-back map
+Since the moduli space $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ parametrizes relative maps to the fibers of $\pi_{S}$, we have a pull-back map
 (2.4) $\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\cong\nu_{*}\omega_{\nu}\ \stackrel{{\scriptstyle f^{*}}}{{\longrightarrow}}\ \mu_{*}\omega_{\mu}\cong\mathbb{E}_{g}$
-over ¯$\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. Since $n>0$, the map (2.4) is injective, so we obtain an exact sequence
+over $\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. Since $n>0$, the map (2.4) is injective, so we obtain an exact sequence
 (2.5) $0\rightarrow\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\rightarrow\mathbb{E}_{g}\rightarrow\mathbb{F}\rightarrow 0\,$,
-where $\mathbb{F}$ is a rank $g-1$ vector bundle on ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$. We therefore have a factorization
-[…]
-on ¯$\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. For the second equality, the restriction of (2.5) to a fiber of $\epsilon_{S}$ is used to obtain
+where $\mathbb{F}$ is a rank g1g-1 vector bundle on $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$. We therefore have a factorization
+(2.6) $\lambda_{g}\ =\ c_{1}(\epsilon_{S}^{*}\big{(}\mathsf{Tan}_{\mathbb{P}^{1}})\big{)}\cdot c_{g-1}(\mathbb{F})\ =\ 2\epsilon_{S}^{*}(\mathsf{pt})\cdot\lambda_{g-1}\,$
+on $\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. For the second equality, the restriction of (2.5) to a fiber of $\epsilon_{S}$ is used to obtain
 $c_{g-1}(\mathbb{F})|_{\epsilon_{S}^{-1}(\mathsf{pt})}=\lambda_{g-1}\,.$
 The factorization (2.6) of $\lambda_{g}$ implies the vanishing
-[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{Tan}_{\mathbb{P}^{1}})=0$
+$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{Tan}_{\mathbb{P}^{1}})=0\$
 by (2.2). ∎
@@
 $\frac{(-1)^{g}(2g-1)!}{24(2g+m-2)!}\left(Q\frac{d}{dQ}\right)^{m-1}E_{2g}(Q)\sum_{i=1}^{m}\int_{\overline{\mathcal{M}}_{g,r}}\prod_{j=1}^{r}\psi_{j}^{k_{j}+1-\delta_{ij}}\cdot\mathsf{F}(\lambda)\lambda_{g}\lambda_{g-1}\,.$
-The integrals over ¯$\overline{\mathcal{M}}_{g,r}$ can be evaluated effectively via Hodge integral techniques.
+The integrals over $\overline{\mathcal{M}}_{g,r}$ can be evaluated effectively via Hodge integral techniques.
 $\bullet$ For $\mathsf{F}=1$, we have the exact evaluation
-¯$\int_{\overline{\mathcal{M}}_{g,r}}\prod_{j=1}^{r}\psi_{j}^{k_{j}+1-\delta_{ij}}\cdot\lambda_{g}\lambda_{g-1}=\frac{|B_{2g}|}{2^{2g-1}(2g)!}\frac{(2g+r-3)!(2k_{i}+1)}{(2k_{1}+1)!!\ldots(2k_{r}+1)!!}\,,$
+$\int_{\overline{\mathcal{M}}_{g,r}}\prod_{j=1}^{r}\psi_{j}^{k_{j}+1-\delta_{ij}}\cdot\lambda_{g}\lambda_{g-1}=\frac{|B_{2g}|}{2^{2g-1}(2g)!}\frac{(2g+r-3)!(2k_{i}+1)}{(2k_{1}+1)!!\ldots(2k_{r}+1)!!}\,$,
 given by the Virasoro constraints of $\mathbb{P}^{2}$ [14]. We arrive at the following:
@@
 There are two connected Gromov-Witten invariants ((i) and (iii)) and two partition counts ((ii) and (iv)) which occur in the connected/disconnect calculus for $\langle(2,1^{n-2})\rangle^{\pi_{\mathbb{C}^{2}},\bullet}$:
-(i) Let ¯$\overline{\mathcal{M}}^{\circ}_{g}(\pi,(2,1^{n-2}))$ be the moduli space of degree $n$ stable relative maps to the fibers of the universal elliptic curve
+(i) Let $\overline{\mathcal{M}}^{\circ}_{g}(\pi,(2,1^{n-2}))$ be the moduli space of degree $n$ stable relative maps to the fibers of the universal elliptic curve
 $\pi:\mathcal{E}\rightarrow\overline{\mathcal{M}}_{1,1}$
@@
 The disconnected invariants $\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}$ can be calculated by degenerating $E$ to a curve of arithmetic genus $1$ with a unique node and applying the correspondence between local Gromov-Witten theory of $\mathbb{P}^{1}$ and quantum cohomology ring of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ [4, 39]:
-(3.2) $-\sum_{g\in\mathbb{Z}}u^{2g-3}\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}=(-i)\cdot\mathsf{trace}\left(\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)\right)=(-i)\cdot\mathsf{Tr}_{n}\cdot(t_{1}+t_{2}),$
-after $-q=e^{iu}$.
+(3.2) $\sum_{g\in\mathbb{Z}}u^{2g-3}\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}=(-i)\cdot\mathsf{trace}\left(\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)\right)=(-i)\cdot\mathsf{Tr}_{n}\cdot(t_{1}+t_{2}),$
+after $q=e^{iu}$.
 (iv) For an integer $l\geq 1$, let $\widetilde{\mathsf{Part}}(l)$ be the count of possibly disconnected unramified covers of $E$ of degree $l$ where each cover is weighted by the order of automorphism group and the number of connected components.
@@
 We then obtain $\widetilde{\mathcal{P}}(x)=\partial_{y}\mathcal{F}(x,y)|_{y=1}=\mathcal{P}(x)\operatorname{log}\mathcal{P}(x)$ by the definition of $\mathsf{Hur}(l,k)$. ∎
-We can now state the main connected/disconnect equation which will play an essential role in the proof of Theorem 1. The proof will be presented in Section 3.4.
+We can now state the main connected/disconnected equation which will play an essential role in the proof of Theorem 1. The proof will be presented in Section 3.4.
 Proposition 10 (Connected/disconnected calculus).
@@
 Lemma 11.
-Under the variable change $-q=e^{iu}$, we have
+Under the variable change $q=e^{iu}$, we have
 $(-i)\sum_{m=1}^{\infty}\mathsf{Tr}_{m}(q)\,Q^{m}=\mathcal{P}(Q)\mathcal{B}(u,Q).$
 Proof.
-The diagonal terms of $\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}$ have been described in [37, Section 2]:
-$\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}=\sum_{r=1}^{\infty}\left(\frac{r}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}-\frac{1}{2}\frac{(-q)+1}{(-q)-1}\right)\alpha_{-r}\alpha_{r}+\ldots,$
+The diagonal terms of $\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}$ have been described in [39, Section 2]:
+$\frac{1}{t_{1}+t_{2}}\mathsf{M}_{D}^{\mathsf{Hilb}^{m}(\mathbb{C}^{2})}=\sum_{r=1}^{\infty}\left(\frac{r}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}-\frac{1}{2}\frac{(-q)+1}{(-q)-1}\right)\alpha_{-r}\alpha_{r}+\ldots,$
 where $\alpha_{-r}$ and $\alpha_{r}$ are the standard creation and annihilation operators on the subspace $\mathcal{F}_{m}\subset\mathcal{F}$ of Fock space
@@
 (3.6) $\begin{split}\frac{-ir}{2}\frac{(-q)^{r}+1}{(-q)^{r}-1}&=\frac{-ir}{2}\frac{e^{iur/2}+e^{-iur/2}}{e^{iur/2}-e^{iur/2}}\\ &=-\frac{r}{2}\cot\left(\frac{ur}{2}\right)\\ &=-\frac{1}{u}+\sum_{h=0}^{\infty}\frac{|B_{2h}|r^{2h}}{(2h)!}u^{2h-1}\,.\end{split}$
-After putting together (3.4), (3.5) and (3.6) and shifting $h=g-1$, we obtain:
+After putting together (3.4), (3.5) and (3.6) and shifting h=g1h=g-1, we obtain:
 $\begin{split}(-i)\sum_{m=1}^{\infty}\mathsf{Tr}_{m}(q)\,Q^{m}&=\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\sum_{r=1}^{\infty}\frac{|B_{2g-2}|}{(2g-2)!}(r^{2g-1}-r)\frac{Q^{r}}{1-Q^{r}}u^{2g-3}\right)\\ &=\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\frac{|B_{2g-2}|}{(2g-2)!}u^{2g-3}\sum_{r=1}^{\infty}(r^{2g-1}-r)(Q^{r}+Q^{2r}+\ldots)\right)\\ &=\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\frac{|B_{2g-2}|}{(2g-2)!}u^{2g-3}\sum_{m=1}^{\infty}Q^{m}\sum_{k\mid m}(k^{2g-1}-k)\right)\\ &=\mathcal{P}(Q)\mathcal{B}(u,Q)\,.\end{split}$
@@
 $\mathcal{P}(Q)\left(\sum_{g=1}^{\infty}\sum_{n=1}^{\infty}\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\circ}u^{2g-3}Q^{n}\right)=i(t_{1}+t_{2})\sum_{n=1}^{\infty}\mathsf{Tr}_{n}(q)\,Q^{n}=-(t_{1}+t_{2})\mathcal{P}(Q)\mathcal{B}(u,Q)\ ,$
-after $-q=e^{iu}$. The connected/disconnected calculus of Proposition 10 then yields
+after $q=e^{iu}$. The connected/disconnected calculus of Proposition 10 then yields
 $\begin{split}\sum_{g=1}^{\infty}\sum_{n=1}^{\infty}\langle(2,1^{n-2})\rangle_{g,n}^{\pi_{\mathbb{C}^{2}},\bullet}Q^{n}u^{2g-3}&=-\frac{1}{24}\frac{t_{1}+t_{2}}{t_{1}t_{2}}(-(t_{1}+t_{2})\mathcal{B}(u,Q))\cdot\widetilde{\mathcal{P}}(Q)+\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\frac{1}{24}\mathcal{B}(u,Q)\mathcal{P}(Q)\\ &=\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\mathcal{B}(u,Q)\mathcal{P}(Q)(1+\log\mathcal{P}(Q))\\ &=\frac{1}{i}\frac{1}{24}\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\left(\sum_{n=1}^{\infty}\mathsf{Tr}_{n}(q)\,Q^{n}\right)\left(1+\sum_{k=1}^{\infty}\sigma_{-1}(k)Q^{k}\right)\,.\end{split}\,$
@@
   • $2\leq m\leq n$,
-  • $\mathbf{k}=(k_{1},\ldots,k_{s})$ is a partition of $n-m$,
-  • $\mathbf{g}=(g_{0},\ldots,g_{s})$ is a partition of $g+s-1$,
+  • $\mathbf{k}=(k_{1},\ldots,k_{s})$ is a partition of nmn-m,
+  • $\mathbf{g}=(g_{0},\ldots,g_{s})$ is a partition of g+s1g+s-1,
 where the parts of $\mathbf{k}$ and $\mathbf{g}$ are positive integers. We define
-¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ =¯$=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{1}}^{\circ}(\pi,(1^{k_{1}}))\times_{\overline{\mathcal{M}}_{1,1}}\ldots\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}})),$
-¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ =¯$=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times\overline{\mathcal{M}}^{\circ}_{g_{1}}(\pi,(1^{k_{1}}))\times\ldots\times\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}}))\,,$
-where ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ is the base change of ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ along the diagonal $\Delta:\overline{\mathcal{M}}_{1,1}\to\overline{\mathcal{M}}_{1,1}\times\ldots\times\overline{\mathcal{M}}_{1,1}$.
-The fixed locus of ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},n,(2,1^{n-2}))$ for the action of the torus $(\mathbb{C}^{*})^{2}$ is the disjoint union of the stacks ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$, up to an automorphism factor which we denote by $\operatorname{Aut}(\mathbf{k},\mathbf{g})$, and the restriction of the virtual class [¯$[\overline{\mathcal{M}}_{g}^{\bullet}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}$ to ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ agrees with $\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}$.
+$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ $=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{1}}^{\circ}(\pi,(1^{k_{1}}))\times_{\overline{\mathcal{M}}_{1,1}}\ldots\times_{\overline{\mathcal{M}}_{1,1}}\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}})),$
+$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ $=\overline{\mathcal{M}}_{g_{0}}^{\circ}(\pi,(2,1^{m-2}))\times\overline{\mathcal{M}}^{\circ}_{g_{1}}(\pi,(1^{k_{1}}))\times\ldots\times\overline{\mathcal{M}}_{g_{s}}^{\circ}(\pi,(1^{k_{s}}))\,$,
+where $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ is the base change of $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ along the diagonal $\Delta:\overline{\mathcal{M}}_{1,1}\to\overline{\mathcal{M}}_{1,1}\times\ldots\times\overline{\mathcal{M}}_{1,1}$.
+The fixed locus of $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},n,(2,1^{n-2}))$ for the action of the torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ is the disjoint union of the stacks $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$, up to an automorphism factor which we denote by $\operatorname{Aut}(\mathbf{k},\mathbf{g})$, and the restriction of the virtual class $\overline{\mathcal{M}}_{g}^{\bullet}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}$ to $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}^{\mathsf{fiber}}$ agrees with $\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}$.
 By the localization formula of [19], we have
-[¯$[\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\frac{\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}}{e_{(\mathbb{C}^{*})^{2}}(N^{vir})}\,,$
-where the contribution of the virtual normal bundle on each factor of ¯$\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ is given by
-(3.8) $\begin{split}\frac{1}{\mathsf{e}_{(\mathbb{C}^{*})^{2}}(N^{vir})}&=\frac{c(\mathbb{E}^{\vee}\otimes t_{1})c(\mathbb{E}^{\vee}\otimes t_{2})}{t_{1}t_{2}}\\ &=-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-1}+\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-2}-(t_{1}+t_{2})\lambda_{g_{i}-1}\lambda_{g_{i}-2}+\ldots\end{split}$
-if $g_{i}>1$, and $-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{1}+1$ if $g_{i}=1$.
+$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},(2,1^{n-2}))]^{vir}=\sum_{(m,\mathbf{k},\mathbf{g})}\frac{1}{\operatorname{Aut}(\mathbf{k},\mathbf{g})}\frac{\Delta^{!}[\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}]^{vir}}{e_{(\mathbb{C}^{*})^{2}}(N^{vir})}\,$,
+where the contribution of the virtual normal bundle on each factor of $\overline{\mathcal{M}}_{m,\mathbf{k},\mathbf{g}}$ is given by
+(3.8) $\begin{split}\frac{1}{\mathsf{e}_{\mathsf{T}}(N^{vir})}&=\frac{c(\mathbb{E}^{\vee}\otimes t_{1})c(\mathbb{E}^{\vee}\otimes t_{2})}{t_{1}t_{2}}\\ &=-\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-1}+\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\lambda_{g_{i}}\lambda_{g_{i}-2}-(t_{1}+t_{2})\lambda_{g_{i}-1}\lambda_{g_{i}-2}+\ldots\end{split}$
+if $g_{i}>1$, and $\frac{t_{1}+t_{2}}{t_{1}t_{2}}\lambda_{1}+1$ if $g_{i}=1$.
 We denote the class (3.8) by $N_{g_{i}}$, and note that the lower order terms in the Hodge classes will not contribute for dimensional reasons.
@@
 where $\operatorname{ev}_{i}:\overline{\mathcal{M}}_{(m,\mathbf{k},\mathbf{g})}\to\overline{\mathcal{M}}_{1,1}$ recovers the isomorphism class of the target curve of the $i$-th map.
-The second equality follows from the decomposition of the diagonal of ¯$\overline{\mathcal{M}}_{1,1}$:
-$[\Delta]=\sum_{j=0}^{s}[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\boxtimes 1\boxtimes[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\in\operatorname{\mathsf{CH}}^{s}(\overline{\mathcal{M}}_{1,1}\times\cdots\times\overline{\mathcal{M}}_{1,1})\,.$
+The second equality follows from the decomposition of the diagonal of $\overline{\mathcal{M}}_{1,1}$:
+$\Delta]=\sum_{j=0}^{s}[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\boxtimes 1\boxtimes[\mathsf{pt}]\boxtimes\ldots\boxtimes[\mathsf{pt}]\in\operatorname{\mathsf{CH}}^{s}(\overline{\mathcal{M}}_{1,1}\times\cdots\times\overline{\mathcal{M}}_{1,1})\,.$
 If $g_{i}>1$, the class $N_{g_{i}}$ contains a $\lambda_{g_{i}}$ factor. Hence, $\langle N_{g_{i}}|(1^{k_{i}})\rangle_{g_{i},k_{i}}^{E,\circ}=0$ unless $g_{i}=1$, in which case
@@
 obtained from the standard degeneration to the normal cone of the section $\mathsf{p}_{1}$.
-We now follow the geometric notation and analysis of Section 2.5. The moduli space of stable maps to the fibers of $\pi$ lies over ¯$\overline{\mathcal{M}}_{1,1}$,
+We now follow the geometric notation and analysis of Section 2.5. The moduli space of stable maps to the fibers of $\pi$ lies over $\overline{\mathcal{M}}_{1,1}$,
 $\epsilon:\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\rightarrow\overline{\mathcal{M}}_{1,1}\,.$
 The universal curve $\mu:\mathcal{C}\rightarrow\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$ carries a universal evaluation map
-$\mathcal{C}$$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$$f$$\mu$$\nu$
+[…]
 to the universally expanded target
-$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\times_{\overline{\mathcal{M}}_{1,1}}\mathcal{E}$¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})\,.$$h$$\nu$$\pi_{M}$
+[…]
 […]
 (3.10) $\epsilon^{*}\mathbb{E}_{1}\cong\nu_{*}\omega_{\mu}\ \stackrel{{\scriptstyle f^{*}}}{{\rightarrow}}\ \mu_{*}\omega_{\mu}\cong\mathbb{E}_{g}$
-over ¯$\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$. Since $k_{i}>0$, the map (3.10) is injective, so we obtain an exact sequence
+over $\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$. Since $k_{i}>0$, the map (3.10) is injective, so we obtain an exact sequence
 $0\rightarrow\epsilon^{*}\mathbb{E}_{1}\rightarrow\mathbb{E}_{g}\rightarrow\mathbb{F}\rightarrow 0\,$,
-where $\mathbb{F}$ is a rank $g-1$ vector bundle on ¯$\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$. We therefore have a factorization
+where $\mathbb{F}$ is a rank g1g-1 vector bundle on $\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})$. We therefore have a factorization
 (3.11) $\lambda_{g}\ =\ c_{1}(\epsilon^{*}\mathbb{E}_{1})\cdot c_{g-1}(\mathbb{F})\ =\ \frac{1}{24}\epsilon^{*}([E])\cdot\lambda_{g-1}\,$
-on ¯$\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$, where $[E]\in\overline{\mathcal{M}}_{1,1}$ is the moduli point of a fixed nonsingular elliptic curve $E$.
+on $\overline{\mathcal{M}}_{g_{i}}(\pi,k_{i})$, where $E]\in\overline{\mathcal{M}}_{1,1}$ is the moduli point of a fixed nonsingular elliptic curve $E$.
 The factorization (3.11) of $\lambda_{g}$ implies
-(3.12) [¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})]^{vir}}\lambda_{g_{i}}\lambda_{g_{i}-1}=\frac{1}{24}\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(E,k_{i})]^{vir}}\lambda^{2}_{g_{i}-1}\,.$
+(3.12) $\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(\pi,k_{i})]^{vir}}\lambda_{g_{i}}\lambda_{g_{i}-1}=\frac{1}{24}\int_{[\overline{\mathcal{M}}^{\circ}_{g_{i}}(E,k_{i})]^{vir}}\lambda^{2}_{g_{i}-1}\,.$
 If $g_{i}>1$, then $\lambda_{g_{i}-1}^{2}=2\lambda_{g_{i}}\lambda_{g_{i}-2}$.
@@
 ∎
-Therefore, the only positive contributions to the right hand side of (3.9) come from triplets $(e,\mathbf{k},\mathbf{g})$ where $\mathbf{g}=(g,1,\ldots,1)$.
+Therefore, the only possible non-zero contributions to the right hand side of (3.9) come from triplets $(e,\mathbf{k},\mathbf{g})$ where $\mathbf{g}=(g,1,\ldots,1)$ with $g\geq 1$.
 $\bullet$ For $g>1$, we have
@@
 $\pi\cup\pi_{P}:\mathcal{E}\cup\mathbb{P}(T_{\mathsf{p_{1}}}\mathcal{E}\oplus\mathbb{C})\longrightarrow\overline{\mathcal{M}}_{1,1},$
-where the gluing identifies the section $\mathsf{p}_{1}$ of $\mathcal{E}$ with the 00-section of $\mathbb{P}(T_{\mathsf{p}_{1}}\mathcal{E}\oplus\mathbb{C})$.
+where the gluing identifies the section $\mathsf{p}_{1}$ of $\mathcal{E}$ with the 0-section of $\mathbb{P}(T_{\mathsf{p}_{1}}\mathcal{E}\oplus\mathbb{C})$.
 The universal section $\mathsf{p}_{1}$ is now the $\infty$-section of $\pi_{P}$.
@@
     Vertex (3.16) vanishes because
-    [¯$[\overline{\mathcal{M}}_{g}(\pi/\mathsf{p}_{1})]^{vir}\cap\operatorname{ev}_{\overline{\mathcal{M}}_{1,1}}^{*}(\mathsf{pt})=[\overline{\mathcal{M}}_{g}(E/p_{1})]^{vir}$
+    $\overline{\mathcal{M}}_{g}(\pi/\mathsf{p}_{1})]^{vir}\cap\operatorname{ev}_{\overline{\mathcal{M}}_{1,1}}^{*}(\mathsf{pt})=[\overline{\mathcal{M}}_{g}(E/p_{1})]^{vir}$
     pairs to zero with $\lambda_{g}$.
   2. $\bullet$
-    Vertex (3.21) vanishes because of the two point conditions over ¯$\overline{\mathcal{M}}_{1,1}$.
+    Vertex (3.21) vanishes because of the two point conditions over $\overline{\mathcal{M}}_{1,1}$.
   3. $\bullet$
@@
 Figure 1. Non-zero contributions to (3.13)
-The vertices (3.18) and (3.20) contribution factors $1$ and $1/2$ respectively (the latter is cancelled by the multiplicity which occurs in the degeneration formula).
+The vertices (3.18) and (3.20) contribution factors 11 and $1/2$ respectively (the latter is cancelled by the multiplicity which occurs in the degeneration formula).
 The above vertex analysis show that (3.13) specializes to:
@@
 Since the integrand has a $\lambda_{g}\lambda_{g-1}$ insertion, there are only two components of the fixed locus that contribute, corresponding to the graphs in Figure 2, see [19].
-[…]
-[…]
+0g1g-1
+g1g-10
 Figure 2. Localization contributions, where half edges correspond to marked points, vertices are contracted components, and edges are non-contracted components.
@@
 To every genus $1$ series $\big{\langle}D^{\star k_{1}},...,D^{\star k_{\ell}}\big{\rangle}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$, there are canonically associated functions
-[…]
+$\big{\{}C_{k,m}\big{\}}_{0\leq k\leq|\mathsf{Part}(n)|-1,\,0\leq m\leq\ell-1}\subset\mathbb{Q}(t_{1},t_{2},q)$
 for which the following equation holds:
@@
 4.2. Proof of Theorem 4
-We adapt the reconstruction strategy of [50, Section 6] to prove the result. Let $m\geq 0$, and fix classes
+We adapt the reconstruction strategy of [52, Section 6] to prove the result. Let $m\geq 0$, and fix classes
 $v_{1},v_{2},v_{3},v_{4},\phi_{1},...,\phi_{m}\in H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))\,.$
-Pulling back Getzler’s relation from ¯$\overline{\mathcal{M}}_{1,4}$ to ¯$\overline{\mathcal{M}}_{1,4+m}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),d)$ and summing over the curve degree $d$, we obtain the following equation (see [26, Equation (27)])
-[…]
-where the relation $\mathsf{X}\sim\mathsf{Y}$ means that the difference $\mathsf{X}-\mathsf{Y}$ is an explicit sum of arbitrary genus 00 invariants and genus $1$ invariants with at most $m+1$ insertions.
+Pulling back Getzler’s relation from $\overline{\mathcal{M}}_{1,4}$ to $\overline{\mathcal{M}}_{1,4+m}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),d)$ and summing over the curve degree $d$, we obtain the following equation (see [27, Equation (27)])
+$3\sum_{h\in S_{4}}\langle v_{h(1)}\star v_{h(2)},v_{h(3)}\star v_{h(4)},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}\,\sim\,4\sum_{h\in S_{4}}\langle v_{h(1)}\star v_{h(2)}\star v_{h(3)},v_{h(4)},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,{\framebox{\tiny{2+m}}}}\,$,
+where the relation $\mathsf{X}\sim\mathsf{Y}$ means that the difference $\mathsf{X}-\mathsf{Y}$ is an explicit sum of arbitrary genus 0 invariants and genus $1$ invariants with at most $m+1$ insertions.
 Getzler’s relation can be equivalently cast in the following form:
@@
 $f(0)\sim 0\,,\quad f(l)\sim 0\,.$
-By expanding the definitions, we see, for $20\leq k\leq l-$ 2,
+By expanding the definitions, we see, for $0\leq k\leq l-2$,
 $-\Psi(a\star D^{*k},b\star D^{*(l-2-k)},D,D)$ $=$ $f(k+2)-2f(k+1)+f(k)-h(l)$
@@
 $f(k+2)-2f(k+1)+f(k)-h(l)\sim 0\,$,
-from (4.2). A linear algebraic result [50, Lemma 6.2] about solutions to the difference equation then yields
+from (4.2). A linear algebraic result [52, Lemma 6.2] about solutions to the difference equation then yields
 (4.4) $f(i)\sim-\frac{i(l-i)}{2}h(l)\,.$
@@
 for all $i,j\geq 0$.
-Repeated use of relation (4.7) shows that $\langle D^{*i},D^{*j},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$ can be expressed in terms of genus 00 invariants and $1$-point genus $1$-functions.
+Repeated use of relation (4.7) shows that $\langle D^{*i},D^{*j},\phi_{1},...,\phi_{m}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{\tiny{2+m}}}$ can be expressed in terms of genus 0 invariants and $1$-point genus $1$-functions.
 Moreover, since we use the divisor equation in the argument, the resulting expressions are linear in $1$-point genus $1$ functions and their $q\frac{d}{dq}$ derivatives.
-By the genus 00 reconstruction result in [37], the genus 00 series can be written as rational functions in matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$.
+By the genus 0 reconstruction result in [39], the genus 0 series can be written as rational functions in matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$.
 The proof of Theorem 4 is complete.
@@
 Since the eigenvectors $v_{i}$ are found by solving the equation $D\star_{\mathsf{t}}v_{i}=e_{i}v_{i}$, the components of $v_{i}$ are rational functions of $e_{i}$ with coefficients in the matrix coefficients of $D\star_{\mathsf{t}}$.
-Therefore, the $\Delta_{i}$ are rational functions in the eigenvalues $e_{i}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$, where the matrix coefficients of $D\star_{\mathsf{t}}$ lie.
-To evaluate the term $\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})$, we need only evaluate symmetric rational functions in the derivatives of the eigenvalues $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$.
-By Proposition 14 of Appendix A, these expressions lie in the field of rational functions of derivatives of the symmetric functions of $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)[[\mathsf{t}]]$.
+Therefore, the $\Delta_{i}$ are rational functions in the eigenvalues $e_{i}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2},q)[[\mathsf{t}]]$, where the matrix coefficients of $D\star_{\mathsf{t}}$ lie.
+To evaluate the term $\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})$, we need only evaluate symmetric rational functions in the derivatives of the eigenvalues $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)[[\mathsf{t}]]$.
+By Proposition 14 of Appendix A, these expressions lie in the field of rational functions of derivatives of the symmetric functions of $e_{i}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)[[\mathsf{t}]]$.
 The outcome is an explicit calculation of
-$\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
+$\nabla_{D^{\star k}}\log(\prod_{i}\Delta_{i})\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2},q)\,.$
 The same argument can be used to calculate derivatives
-$\nabla_{v_{1}}\nabla_{v_{2}}\cdots\nabla_{v_{k}}\log(\prod_{i}\Delta_{i})\,\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
+$\nabla_{v_{1}}\nabla_{v_{2}}\cdots\nabla_{v_{k}}\log(\prod_{i}\Delta_{i})\,\big{|}_{\mathsf{t}=0}\in\mathbb{Q}(t_{1},t_{2},q)\,.$
 Next, we consider the term $\nabla_{D^{\star k}}u_{i}\,|_{\mathsf{t}=0}$ of (4.8).
@@
 $\delta_{l}=\langle D,..,D\rangle_{1,\,\framebox{\tiny{$\ell$}}}-\left(q\frac{d}{dq}\right)^{l-1}\frac{1}{48}\nabla_{D}\log\Big{(}\,\prod_{i=1}^{|\mathsf{Part(n)}|}\Delta_{i}\Big{)}\big{|}_{\mathsf{t}=0}.$
-By Theorem 1 and the discussion in Section 4.3, $\delta_{l}\in\mathbb{Q}(t_{1},t_{2})(q)$ can be explicitly calculated.
+By Theorem 1 and the discussion in Section 4.3, $\delta_{l}\in\mathbb{Q}(t_{1},t_{2},q)$ can be explicitly calculated.
 By (4.8), we have, for $\ell\geq 1$,
@@
 By the construction of the $\mathsf{R}$-matrix [15], the derivatives $q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$ are given by
-$q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}=\sum_{l}\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}\cdot\left(\nabla_{D}u_{l}\,\big{|}_{\mathsf{t}=0}-\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\right)\cdot\mathsf{R}_{li}^{1}\,\big{|}_{\mathsf{t}=0}\,,$
-where the off-diagonal terms $\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}$ are computed by the equation
+$q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}=\sum_{l}\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}\cdot\left(\nabla_{D}u_{l}\,\big{|}_{\mathsf{t}=0}-\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\right)\cdot\mathsf{R}_{li}^{1}\,\big{|}_{\mathsf{t}=0}\,.$
+The off-diagonal terms $\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}$ are computed by the equation
 $\Psi^{-1}\cdot q\frac{d}{dq}\Psi=[\nabla_{D}U,\mathsf{R}^{1}]\,$,
 where $\nabla_{D}U$ is the diagonal matrix with diagonal entries $\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}$ and $\Psi$ is the matrix whose columns are normalized eigenvectors.
-The functions $q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$ are therefore rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2})(q)$, where the matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$ lie.
+The functions $q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$ are therefore rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2},q)$, where the matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$ lie.
 The higher order derivatives
 $\left(q\frac{d}{dq}\right)^{k>1}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$
-are rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$.
+are rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$.
 We would like to calculate the diagonal terms $\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$. Define the column vector
@@
 Here, $\mathsf{A}$ is a $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|^{2}$ matrix of the shape
-$\mathsf{A}=(\mathsf{W}\,|\,\widehat{\mathsf{W}}),$
+[…]
 where $\mathsf{W}$ is the $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|$ matrix given by
@@
 The matrix $\mathsf{W}$ is the Wronskian of the eigenvalues $e_{i}\,|_{\mathsf{t}=0}$ as functions of $\log(q)$.
-Since the eigenvalues are analytic functions of $\log(q)$, a result of Boecher (see [49, Lemma 1.12]) implies that the eigenvalues are linearly independent over $\mathbb{Q}(t_{1},t_{2})$ if and only if $\det(\mathsf{W})$ is not identically 00.
+Since the eigenvalues are analytic functions of $\log(q)$, a result of Boecher (see [51, Lemma 1.12]) implies that the eigenvalues are linearly independent over $\mathbb{Q}(t_{1},t_{2})$ if and only if $\det(\mathsf{W})$ is not identically 0.
 If $\det(\mathsf{W})$ is nonzero, then $\mathsf{W}$ is invertible. Equation (4.10) then implies
+(4.11) $\mathsf{W}^{-1}\,2\vec{\delta}=\mathsf{W}^{-1}\mathsf{A}\,\vec{r}.$
+The $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|^{2}$ matrix $\mathsf{W}^{-1}\mathsf{A}$ is of the form
 […]
-The $|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|^{2}$ matrix $\mathsf{W}^{-1}\mathsf{A}$ is of the form
-$\mathsf{W}^{-1}\mathsf{A}=(\mathsf{I}_{|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|}\,|\,\mathsf{W}^{-1}\widehat{\mathsf{W}})\,.$
 Via the identity matrix $\mathsf{I}_{|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|}$, equation (4.11) yields equations for
@@
 $\Big{\{}\,\langle D,...,D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1,\,\framebox{{\tiny$\ell$}}}=\left(q\frac{d}{dq}\right)^{\ell-1}\langle D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}\,\Big{\}}_{k=1}^{|\mathsf{Part}(n)|}$
-and rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$. ∎
+and rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$. ∎
 4.5. Proof of Theorem 6
@@
 $e_{1}\,\big{|}_{\mathsf{t}=0}\,,\,\ldots\,,\,e_{\mathsf{Part}(n)}\,\big{|}_{\mathsf{t}=0}$
-with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$ are equivariant under permutations of the indices. Therefore, after substitution in
+with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$ are equivariant under permutations of the indices. Therefore, after substitution in
 $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}=\frac{1}{2}\sum_{i=1}^{|\mathsf{Part}(n)|}\mathsf{R}^{1}_{ii}\,\big{|}_{\mathsf{t}=0}\cdot\nabla_{D^{\star k}}u_{i}\,\big{|}_{\mathsf{t}=0}+\frac{1}{48}\nabla_{D^{\star k}}\log\Big{(}\,\prod_{i=1}^{|\mathsf{Part(n)|}}\Delta_{i}\Big{)}\,\big{|}_{\mathsf{t}=0}\,$,
-the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ are symmetric rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$.
+the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ are symmetric rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$.
 After an application of Proposition 14 of Appendix A, the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ can be effectively reconstructed from $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ and $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)$.
@@
 For $2\leq n\leq 5$ and every partition $\mu$ of $n$, we present here closed formulas for the 1-point series
-$\langle\,\mu\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
+$\langle\,\mu\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2},q)\,.$
 We use a combination of inputs: the full genus 0 theory, the evaluations (5.1) and (5.2), and Getzler’s equation.
@@
 $\bullet$ For $n=4$, we have:
-$\begin{split}&\langle(1,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{109}{12}\,,\\[5.0pt] &\langle\,(2,1,1)\,\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{35q^{5}-28q^{4}+23q^{3}+23q^{2}-28q+35}{2(q-1)(q^{2}+1)(q^{2}-q+1)}\,,\end{split}$
-$\begin{split}\langle(2,2)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot&\frac{3}{16(q^{2}+1)^{2}(q-1)^{2}}\cdot\\ &\ \left((t_{1}^{2}+\frac{1}{18}t_{1}t_{2}+t_{2}^{2})(q^{6}+1)+(\frac{2}{9}t_{1}^{2}+5t_{1}t_{2}+\frac{2}{9}t_{2}^{2})(q^{5}+q)\right.\\ &\left.+(\frac{11}{9}t_{1}^{2}-\frac{77}{18}t_{1}t_{2}+\frac{11}{9}t_{2}^{2})(q^{4}+q^{2})+(4t_{1}^{2}+\frac{170}{9}t_{1}t_{2}+4t_{2}^{2})q^{3}\right)\,,\end{split}$
-$\begin{split}\langle(3,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot&\frac{1}{2(q^{2}-q+1)^{2}(q^{2}+1)^{2}}\cdot\\[5.0pt] &\left((t_{1}^{2}+\frac{23}{36}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{6}t_{1}^{2}-\frac{41}{36}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{7}+q)+(\frac{1}{3}t_{1}^{2}-\frac{257}{36}t_{1}t_{2}+\frac{1}{3}t_{2}^{2})(q^{6}+q^{2})\right.\\ &\left.+(\frac{5}{6}t_{1}^{2}+\frac{137}{12}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{5}+q^{3})-(\frac{8}{3}t_{1}^{2}+\frac{170}{6}t_{1}t_{2}+\frac{8}{3}t_{2}^{2})q^{4}\right)\,,\\[5.0pt] \langle(4)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot&\frac{(q+1)}{4(q^{2}-q+1)(q^{2}+1)^{3}(q-1)}\cdot\\[5.0pt] &\left((t_{1}^{2}-\frac{5}{6}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{3}t_{1}^{2}-\frac{25}{6}t_{1}t_{2}+\frac{5}{3}t_{2}^{2})(q^{7}+q)+(2t_{1}^{2}-13t_{1}t_{2}+2t_{2}^{2})(q^{6}+q^{2})\right.\\[5.0pt] &\left.+(3t_{1}^{2}+\frac{69}{2}t_{1}t_{2}+3t_{2}^{2})(q^{5}+q^{3})-(\frac{10}{3}t_{1}^{2}+39t_{1}t_{2}+\frac{10}{3}t_{2}^{2})q^{4}\right)\,.\end{split}$
+$\langle(1,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{109}{12}\,$,
+$\langle\,(2,1,1)\,\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{35q^{5}-28q^{4}+23q^{3}+23q^{2}-28q+35}{2(q-1)(q^{2}+1)(q^{2}-q+1)}\,$,
+$\hskip-20.0pt\langle(2,2)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot\frac{3}{16(q^{2}+1)^{2}(q-1)^{2}}\cdot$
+$\ \left((t_{1}^{2}+\frac{1}{18}t_{1}t_{2}+t_{2}^{2})(q^{6}+1)+(\frac{2}{9}t_{1}^{2}+5t_{1}t_{2}+\frac{2}{9}t_{2}^{2})(q^{5}+q)\right$.
+$\left.+(\frac{11}{9}t_{1}^{2}-\frac{77}{18}t_{1}t_{2}+\frac{11}{9}t_{2}^{2})(q^{4}+q^{2})+(4t_{1}^{2}+\frac{170}{9}t_{1}t_{2}+4t_{2}^{2})q^{3}\right)\,$,
+$\langle(3,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot\frac{1}{2(q^{2}-q+1)^{2}(q^{2}+1)^{2}}\cdot$
+$\left((t_{1}^{2}+\frac{23}{36}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{6}t_{1}^{2}-\frac{41}{36}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{7}+q)+(\frac{1}{3}t_{1}^{2}-\frac{257}{36}t_{1}t_{2}+\frac{1}{3}t_{2}^{2})(q^{6}+q^{2})\right$.
+$\left.+(\frac{5}{6}t_{1}^{2}+\frac{137}{12}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{5}+q^{3})-(\frac{8}{3}t_{1}^{2}+\frac{170}{6}t_{1}t_{2}+\frac{8}{3}t_{2}^{2})q^{4}\right)\,$,
+$\langle(4)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{(q+1)}{4(q^{2}-q+1)(q^{2}+1)^{3}(q-1)}\cdot$
+$\left((t_{1}^{2}-\frac{5}{6}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{3}t_{1}^{2}-\frac{25}{6}t_{1}t_{2}+\frac{5}{3}t_{2}^{2})(q^{7}+q)+(2t_{1}^{2}-13t_{1}t_{2}+2t_{2}^{2})(q^{6}+q^{2})\right$.
+$\left.+(3t_{1}^{2}+\frac{69}{2}t_{1}t_{2}+3t_{2}^{2})(q^{5}+q^{3})-(\frac{10}{3}t_{1}^{2}+39t_{1}t_{2}+\frac{10}{3}t_{2}^{2})q^{4}\right)\,.$
 $\bullet$ For $n=5$, we have:
@@
 We will write remaining series in terms of the traces [^25] The subscript $m$ of $\mathsf{Tr}_{m}^{\mu}$ is redundant since $m=|\mu|$, but is included for clarity. of quantum multiplication,
-$\mathsf{Tr}_{m}^{\mu}=\mathsf{trace}\left(\,\mu\,\star:QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\to QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\,\right)\,.$
+$\mathsf{Tr}_{m}^{\mu}=\mathsf{trace}\left(\,|\mu\rangle\,\star:QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\to QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\,\right)\,.$
 The trace which appears in Theorem 1 can be written as
@@
 Then, we have:
-$\begin{split}\langle(2,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=-\frac{t_{1}+t_{2}}{18t_{1}t_{2}}\mathsf{Tr}_{2}^{(2)}-\frac{t_{1}+t_{2}}{16t_{1}t_{2}}\mathsf{Tr}_{3}^{(2,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{4}^{(2,1,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{5}^{(2,1,1,1)}\\ &=\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(\frac{272q^{9}-539q^{8}+760q^{7}-629q^{6}+302q^{5}+302q^{4}-629q^{3}+760q^{2}-539q+272}{6(q-1)(q^{2}+1)(q^{2}-q+1)(q^{4}-q^{3}+q^{2}-q+1)}\right)}}\,,\end{split}$
-$\begin{split}\langle(2,2,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{775t_{1}^{2}+733t_{1}t_{2}+775t_{2}^{2}}{1200(t_{1}+t_{2})}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-10t_{1}^{2}-13t_{1}t_{2}-10t_{2}^{2}}{240t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-3}{200(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}+\frac{-25t_{1}^{2}-68t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}\\[5.0pt] &+\frac{50t_{1}^{2}-139t_{1}t_{2}+50t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{-25t_{1}^{2}-44t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,,\end{split}$
-$\begin{split}\langle(3,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{175t_{1}^{2}+77t_{1}t_{2}+175t_{2}^{2}}{300(t_{1}+t_{2})}+\frac{-225t_{1}^{2}-532t_{1}t_{2}-225t_{2}^{2}}{3600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{1}{120(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}\\[5.0pt] &+\frac{-3}{100(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{-25t_{1}^{2}-47t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-2}{75(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}\\[5.0pt] &+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,,\end{split}$
-$\begin{split}\langle(3,2)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{865t_{1}^{2}+1556t_{1}t_{2}+865t_{2}^{2}}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-158t_{1}^{2}-91t_{1}t_{2}-158t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{20t_{1}^{2}+79t_{1}t_{2}+20t_{2}^{2}}{450(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}+\frac{-1}{300(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}\\[5.0pt] &+\frac{35t_{1}^{2}-77t_{1}t_{2}+35t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-25t_{1}^{2}-77t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{1}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}+\frac{-25t_{1}^{2}-39t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}\\[5.0pt] &+\frac{-23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}+\frac{-4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,,\end{split}$
-$\begin{split}\langle(4,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{-430t_{1}^{2}+743t_{1}t_{2}-430t_{2}^{2}}{2400(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-461t_{1}^{2}+503t_{1}t_{2}-461t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{925t_{1}^{2}+1771t_{1}t_{2}+925t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}\\[5.0pt] &+\frac{-100t_{1}^{2}-223t_{1}t_{2}-100t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}+\frac{-113t_{1}t_{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-1}{50(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{-25t_{1}^{2}-63t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}\\[5.0pt] &+\frac{1}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}+\frac{-11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,,\end{split}$
-$\begin{split}\langle(5)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}&=\frac{t_{1}t_{2}(110t_{1}^{2}-1391t_{1}t_{2}+110t_{2}^{2})}{600(t_{1}+t_{2})}+\frac{-75t_{1}^{2}-53t_{1}t_{2}-75t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-50t_{1}^{2}+77t_{1}t_{2}-50t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}\\[5.0pt] &+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{75t_{1}^{2}+53t_{1}t_{2}+75t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-5t_{1}^{2}-16t_{1}t_{2}-5t_{2}^{2}}{120t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(5)}\\[5.0pt] &+\frac{175t_{1}^{2}+94t_{1}t_{2}+175t_{2}^{2}}{150(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{44t_{1}^{2}+45t_{1}t_{2}+44t_{2}^{2}}{120(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{-60t_{1}^{2}-53t_{1}t_{2}-60t_{2}^{2}}{200(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,1,1)}\\[5.0pt] &+\frac{15t_{1}^{2}+23t_{1}t_{2}+15t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,.\end{split}$
+$\langle(2,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,-\frac{t_{1}+t_{2}}{18t_{1}t_{2}}\mathsf{Tr}_{2}^{(2)}-\frac{t_{1}+t_{2}}{16t_{1}t_{2}}\mathsf{Tr}_{3}^{(2,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{4}^{(2,1,1)}-\frac{t_{1}+t_{2}}{24t_{1}t_{2}}\mathsf{Tr}_{5}^{(2,1,1,1)}$
+$=$ $\,\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot{{\left(\frac{272q^{9}-539q^{8}+760q^{7}-629q^{6}+302q^{5}+302q^{4}-629q^{3}+760q^{2}-539q+272}{6(q-1)(q^{2}+1)(q^{2}-q+1)(q^{4}-q^{3}+q^{2}-q+1)}\right)}}\,$,
+$\langle(2,2,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{775t_{1}^{2}+733t_{1}t_{2}+775t_{2}^{2}}{1200(t_{1}+t_{2})}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-10t_{1}^{2}-13t_{1}t_{2}-10t_{2}^{2}}{240t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}$
+$+\frac{-3}{200(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}+\frac{-25t_{1}^{2}-68t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{1}{200(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}$
+$+\frac{50t_{1}^{2}-139t_{1}t_{2}+50t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{-25t_{1}^{2}-44t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,$,
+$\langle(3,1,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{175t_{1}^{2}+77t_{1}t_{2}+175t_{2}^{2}}{300(t_{1}+t_{2})}+\frac{-225t_{1}^{2}-532t_{1}t_{2}-225t_{2}^{2}}{3600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{1}{120(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(3,1)}$
+$+\frac{-3}{100(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{-25t_{1}^{2}-47t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-2}{75(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}$
+$+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{1}{100(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,$,
+$\langle(3,2)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{865t_{1}^{2}+1556t_{1}t_{2}+865t_{2}^{2}}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-158t_{1}^{2}-91t_{1}t_{2}-158t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{20t_{1}^{2}+79t_{1}t_{2}+20t_{2}^{2}}{450(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}+\frac{-1}{300(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}$
+$+\frac{35t_{1}^{2}-77t_{1}t_{2}+35t_{2}^{2}}{900(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-25t_{1}^{2}-77t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{1}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}+\frac{-25t_{1}^{2}-39t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}$
+$+\frac{-23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}+\frac{-4}{225(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{23}{1800(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,$,
+$\langle(4,1)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{-430t_{1}^{2}+743t_{1}t_{2}-430t_{2}^{2}}{2400(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}+\frac{-461t_{1}^{2}+503t_{1}t_{2}-461t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(2,1)}+\frac{925t_{1}^{2}+1771t_{1}t_{2}+925t_{2}^{2}}{4800(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,1,1)}$
+$+\frac{-100t_{1}^{2}-223t_{1}t_{2}-100t_{2}^{2}}{2400t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(4)}+\frac{-113t_{1}t_{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,1,1,1)}+\frac{-1}{50(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,2)}+\frac{-25t_{1}^{2}-63t_{1}t_{2}-25t_{2}^{2}}{600t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(4,1)}$
+$+\frac{1}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(3)}+\frac{-11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(3,1)}+\frac{7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,2)}$
+$+\frac{-7}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,2,1)}+\frac{11}{600(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(3,1,1)}\,$,
+$\langle(5)\rangle_{1}^{\mathsf{Hilb}^{5}(\mathbb{C}^{2})}=$ $\,\frac{t_{1}t_{2}(110t_{1}^{2}-1391t_{1}t_{2}+110t_{2}^{2})}{600(t_{1}+t_{2})}+\frac{-75t_{1}^{2}-53t_{1}t_{2}-75t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{3}^{(3)}+\frac{-50t_{1}^{2}+77t_{1}t_{2}-50t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{4}^{(2,2)}$
+$+\frac{-25t_{1}^{2}-59t_{1}t_{2}-25t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(2,2,1)}+\frac{75t_{1}^{2}+53t_{1}t_{2}+75t_{2}^{2}}{600(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(3,1,1)}+\frac{-5t_{1}^{2}-16t_{1}t_{2}-5t_{2}^{2}}{120t_{1}t_{2}(t_{1}+t_{2})}\mathsf{Tr}_{5}^{(5)}$
+$+\frac{175t_{1}^{2}+94t_{1}t_{2}+175t_{2}^{2}}{150(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{2}^{(2)}+\frac{44t_{1}^{2}+45t_{1}t_{2}+44t_{2}^{2}}{120(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{3}^{(2,1)}+\frac{-60t_{1}^{2}-53t_{1}t_{2}-60t_{2}^{2}}{200(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{4}^{(2,1,1)}$
+$+\frac{15t_{1}^{2}+23t_{1}t_{2}+15t_{2}^{2}}{300(t_{1}+t_{2})}\mathsf{Tr}_{2}^{(2)}\cdot\mathsf{Tr}_{5}^{(2,1,1,1)}\,.$
 $\bullet$ There is no obstruction (apart from expected nondegeneracies) to extending the above tables to higher $n$.
@@
 Proposition 14.
-$\mathsf{SymDf}\subseteq\mathbb{K}[\mathsf{Ds}][1/\Delta]$
+$\mathsf{SymDf}\subseteq\mathbb{K}[\mathsf{Ds}][1/\Delta]$ .
 Proof.